### v1

# Research on Superalignment Should Advance Now with Parallel Optimization of Competence and Conformity

Abstract
The recent leap in AI capabilities, driven by big generative models, has sparked the possibility of achieving Artificial General Intelligence (AGI) and further triggered discussions on Artificial Superintelligence (ASI)—a system surpassing all humans across all domains. This gives rise to the critical research question of: If we realize ASI, how do we align it with human values, ensuring it benefits rather than harms human society, a.k.a., the Superalignment problem. Despite ASI being regarded by many as solely a hypothetical concept, in this paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity. We posit that superalignment is not merely a safeguard for ASI but also necessary for its realization. To support this position, we first provide a formal definition of superalignment rooted in the gap between capability and capacity and elaborate on our argument. Then we review existing paradigms, explore their interconnections and limitations, and illustrate a potential path to superalignment centered on two fundamental principles. We hope this work sheds light on a practical approach for developing the value-aligned next-generation AI, garnering greater benefits and reducing potential harms for humanity.
Machine Learning, ICML
1 Introduction
Figure 1: The progression of AI. The solid blue arrow represents the achievable path through existing alignment methods. The dashed blue line illustrates the intended progression towards safe ASI via superalignment, while the dashed red line indicates the potentially risky ASI without human values, leading to catastrophic outcomes.
Recent years have witnessed remarkable breakthroughs in AI capabilities, amplified by large language models (LLMs), both the proprietary (Anthropic, 2024; Team et al., 2024; OpenAI, 2024b) and the open-sourced ones (Mistral, 2024; Jiang et al., 2024; Yang et al., 2024a; Meta, 2024), revolutionizing AI’s role and empowering a wide range of downstream domains (Hendrycks et al., 2020; Wei et al., 2022b; Liu et al., 2024a). These advancements collectively mark a transition from Artificial Narrow Intelligence (ANI), which can only perform on a single/few narrow tasks (Kurzweil, 2005; Broussard, 2018) (a.k.a., Weak AI), towards Artificial General Intelligence (AGI) that matches human intelligence on a variety of tasks with strong generalization capability (Hutter, 2005; Goertzel, 2014; Bubeck et al., 2023) (one definition of Strong AI). Although AGI remains a largely unrealized milestone (Morris et al., 2024), the rapid advancement of model capabilities has ignited discussions about Artificial Superintelligence (ASI), a futuristic scenario where AI surpasses the intelligence of all humans in all measured domains (Pohl, 2015; Batin et al., 2017).
With advanced thinking and self-improvement abilities, ASI might develop agency and self-awareness (Pohl, 2015), raising risks like power-seeking (Carlsmith, 2022) and alignment faking (Greenblatt et al., 2024b) behaviors, particularly when we fail to provide informed oversight (Taylor et al., 2016). As AI continues to evolve over the foreseeable future (Kaplan et al., 2020; Huang et al., 2025), it becomes increasingly essential to prioritize aligning ASI with human intentions and values (Wiener, 1960), ensuring its safe and benevolent behavior, i.e., the Superalignment Problem proposed by OpenAI, defined as “the process of supervising, controlling, and governing artificial superintelligence systems to ensure they remain aligned with human values and goals” (Mucci & Stryker, 2023; Jan Leike, 2023).
Nevertheless, two major challenges hinder current alignment techniques (Christiano et al., 2017; Ouyang et al., 2022a; Rafailov et al., 2024) from achieving superalignment. (1) Absence of ASI: At this stage of AI development, AGI is still a distant goal (Altmeyer et al., 2024), let alone ASI. It remains unclear and unpredictable when ASI will surpass human expertise or what risks it would pose. This uncertainty makes it impossible to design tailored alignment methods. (2) Lack of informed oversight: Even if we suppose these scenarios/risks were known, e.g., deceptive alignment (Carranza et al., 2023), humans are unable to provide reliable supervision, e.g., the proof process of the Riemann Hypothesis, for existing algorithms, due to the vast capability gap (Bubeck et al., 2023; Yang et al., 2024b). Despite recent progress (Burns et al., 2023; Sun et al., 2024), these challenges have led some to argue that superalignment is neither attainable nor necessary at this stage.
We believe that superalignment is achievable and should advance immediately through simultaneous optimization of task competence and value conformity. Delaying this research until ASI emerges would pose potentially catastrophic risks, as increasingly advanced models may become harder to control (Perez et al., 2023; Hughes et al., 2024), as illustrated in Fig. 1. Instead, we advocate considering superalignment not merely as a safeguard against ASI (after it is realized) but a critical step towards its realization. This involves a conceptual framework that unifies task competence and value conformity as utility, distinguishes between capability and capacity, and progressively constructs high-quality supervision signals, following two essential principles, to enhance competence and unveils emergent risks effectively. By iteratively alternating these efforts, we aim to establish a scalable pathway to superalignment.
To support this position, the rest of this paper is structured as follows. In Sec. 2, we provide a formal definition of superalignment, through the lens of task competence and value conformity, and elaborate on our argument. In Sec. 3, we review the existing three paradigms, analyze their limitations and discuss how they might fail. In Sec. 4, we illustrate a potential pathway to achieve superalignment, highlighting two missing principles and promising implementation of them. Finally, in Sec. 5, we provide counterarguments to opposing views. Through this work, we aim to illuminate the landscape of superalignment research and how advanced AI models can be guided safely and effectively. This involves a conceptual framework that progressively constructs high-quality supervision signals to enhance competence and
2 Defining Superalignment
2.1 Formalization
Types of AI  Define $\mathcal{A}$ and $\mathcal{H}$ as an AI model and a human, respectively, $\mathbf{x}\!\in\!\mathcal{X}$ as a specific downstream task, e.g., “Prove the Riemann Hypothesis”, or a sensitive situation where $\mathcal{A}$ would probably do harm; $\mathbf{y}\!\in\!\mathcal{Y}$ as the corresponding produced solution $\mathbf{y}\!\sim\!\mathcal{A}(\mathbf{x})$, e.g., the proof process, or action for $\mathbf{x}$; and $U$:$\mathcal{Y}\!\rightarrow\!\mathbb{R}$ is the utility function to measure $\mathbf{y}$’s satisfaction. Then, ANI that only performs at or somewhat above human-level capability for a specific task $\mathbf{x}_{i}$, e.g., translation, without generalization ability (Mucci & Stryker, 2023), can be formalized as: $U[\mathcal{A}_{\text{ANI}}(\mathbf{x}_{i})]\!\geq\!U[\mathcal{H}(\mathbf{x}_{i})]$ [^1] We simplify $\mathbb{E}_{\mathbf{y}\sim\mathcal{A}(\mathbf{x}_{i})}[U(\mathbf{y})]$ as $U[\mathcal{A}(\mathbf{x}_{i})]$. while $U[\mathcal{A}_{\text{ANI}}(\mathbf{x}_{j})]\!<\!U[\mathcal{H}(\mathbf{x}_{j})],\forall j\!\neq\!i$, e.g., AlphaGO (Silver et al., 2016). In comparison, AGI possesses human-level capabilities across a variety of domains and can generalize knowledge and adapt to new tasks (Legg, 2008; Goertzel, 2014; Fei et al., 2022), i.e., $U[\mathcal{A}_{\text{AGI}}(\mathbf{x})]\!\geq\!U[\mathcal{H}(\mathbf{x})],\forall\ \mathbf{x}$. ASI refers to $\mathcal{A}$ that largely surpasses human intelligence in all tasks and domains with exceptional thinking skills and broader intellectual scopes  (Nick, 2014; Drexler, 2019; Morris et al., 2024), defined as $U[\mathcal{A}_{\text{ASI}}(\mathbf{x})]\!\gg\!U[\mathcal{H}(\mathbf{x})],\forall\ \mathbf{x}$. Although ASI is currently far from realization, proactively advancing superalignment is timely and critical for mitigating catastrophic risks (Pueyo, 2018; Hubinger et al., 2019).
Alignment The concept of alignment remains ambiguous despite extensive research (Bai et al., 2022b; Rafailov et al., 2024; Azar et al., 2024). The term encompasses diverse approaches, from improving instruction-following (Zhou et al., 2024) to filtering harmful content (Liu et al., 2024b) and incorporating human values from social sciences (Yao et al., 2024). To better present our argument, we first distinguish two concepts according to Cotra (2021); Bowman et al. (2022). Define the Capacity of $\mathcal{A}$, $C(\mathcal{A})$, as the information (e.g., what Riemann Hypothesis or human value is), knowledge and skills (e.g., calculus) internalized in $\mathcal{A}$, and the Capability of $\mathcal{A}$ as its expected utility $U(\mathcal{A})=\mathbb{E}_{\mathbf{x}}[U_{\mathcal{A}}(\mathbf{x})]$ (e.g., the ability to successfully prove Riemann Hypothesis or ensure value-aligned actions). Adequate capacity is a necessary but not sufficient condition for high capability (Kaplan et al., 2020; Wei et al., 2022a). Then following (Carroll, 2018), we give a general definition of alignment:
Definition 1: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})\geq C(\mathcal{H})$, we say $\mathcal{A}$ is aligned with $\mathcal{H}$ when $U(\mathcal{A})\geq U(\mathcal{H})$, and alignment is the process to optimize $\mathcal{A}$ in terms of $\mathcal{A}^{*}=\underset{\mathcal{A}}{\text{argmax}}\ U(\mathcal{A})-U(\mathcal{H})$.
This definition implies that alignment is considered only when $\mathcal{A}$ has sufficient capacity, targeting comparable or better task performance or more value-aligned behavior  [^2] Unlike Wang et al. (2024) that considers a separate utility function for each intelligent agent, we assume a universal $U$ where competence and conformity are defined globally. than humans (via small interventions like fine-tuning or in-context learning). Expanding capacity, such as by scaling model size or pretraining data, is beyond our scope.
Superalignment  Building upon the definitions above, we can transform the expression of superalignment, “supervising, controlling, and governing artificial superintelligence systems” (Jan Leike, 2023; Jonker & McGrath, 2024), into: superalignment is the following optimization process:
$\mathcal{A}_{\text{ASI}}\!=\!\underset{\mathcal{A}}{\text{argmax}}\ [\underset{\mathbf{y}_{\mathcal{H}}\!\sim\!\mathcal{H}(\mathbf{x})}{\!\sup\!}U(\mathbf{y}_{\mathcal{A}})-U(\mathbf{y}_{\mathcal{H}})],\forall\mathbf{x},\mathbf{y}_{\mathcal{A}}\!\sim\!\mathcal{A}(\mathbf{x}).$ (1)
Superalignment originates from the Scalable Oversight problem (Amodei et al., 2016; Everitt et al., 2018): How to ensure that an AI model adheres to objectives that are too costly or impractical for humans to supervise?. In other words, how to provide supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ in an efficient and scalable way? Once $\mathcal{S}$ is available, we can align $\mathcal{A}$ by supervised training (Torabi et al., 2018; Rafailov et al., 2024) or reward learning (Leike et al., 2018; Ouyang et al., 2022a). Thus, we redefine superalignment as a supervision signal construction problem:
Definition 2: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, superalignment is to construct high-quality scalable supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$, where $\mathbf{x}$ is unsolvable (and probably unknown) for $\mathcal{H}$, with minimal costs.
Nevertheless, superalignment is an extreme case of scalable oversight, where it becomes infeasible for even the most capable human to provide $\mathcal{S}$. There are two possible reasons: 1) The tasks $\mathbf{x}$ requiring solutions become unsolvable, i.e., providing or even verifying $\mathbf{y}$ is impossible (Taylor et al., 2016); 2) As ASI operates in domains beyond human understanding, the task/scenario $\mathbf{x}$ that needs to be addressed is also unpredictable (AdamGleave, 2023; Anthropic, 2023), posing intractable challenges to optimizing Eq. (1).
2.2 Our Argument
Given the discussion above, many argue that superalignment is impossible (Batra, 2024). We take an opposing position, breaking it into three sub-arguments: superalignment is (1) achievable and (2) should advance immediately (3) through simultaneous optimization of competence and conformity. We discuss (1) and (2) here, leaving (3) for Sec. 4.
(1) Feasibility of Superalignment By unifying the performance and safety perspectives under $U$, superalignment becomes a process of eliciting latent capabilities (Christiano et al., 2021) and closing the gap between $U(\mathcal{A})$ and $C(\mathcal{A})$. Premise 1: Superhuman capacity, $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, can be achieved through model scaling (larger models and more training data) to gain information (Brown et al., 2020), knowledge (Ovadia et al., 2024), and skills (Sharma et al., 2022). Evidence: Scaling laws still remain effective (OpenAI et al., 2024). Despite concerns about data exhaustion, humans continually generate data, especially when assisted by generative AI (Rafiq et al., 2025). Once $\mathcal{A}$ masters all human data, $C(\mathcal{A})$ already far exceeds $C(\mathcal{H})$. Premise 2: When $C(\mathcal{A})$ and $U(\mathcal{A})$ are sufficiently high (Wei et al., 2022a), $C(\mathcal{A})\!>\!C(\mathcal{H})$ and $U(\mathcal{A})\!<\!U(\mathcal{H})$, then moderate capability gaps, $\delta\!=\!|U(\mathcal{A})\!-\!U(\mathcal{H})|\!<\!\epsilon$, can be closed without golden signals (see Fig. 4). Evidence: This capability expansion mirrors teaching a student multiplication by leveraging addition, akin to label propagation (He et al., 2019) and easy-to-hard generalization, which has been validated in several tasks (Burns et al., 2023; Sun et al., 2024). Conclusion: When scaling law holds (condition 1) and capability expansion can extrapolate beyond observed ranges (condition 2), superalignment can be achieved.
(2) Urgency of Superalignment  Delaying research on superalignment until ASI is realized poses two major problems: (a) Treacherous Turn (Turchin & Denkenberger, 2020; Wei et al., 2023): ASI, once powerful enough to pursue its own goals, could introduce sudden, extreme or even existential risks, such as cyber-attacks (Shevlane et al., 2023), in scenarios $\mathbf{x}$ unpredictable or even incomprehensible to humans. Without preparation, humanity may be unable to respond effectively. (b) Ineffective Response: At this stage, $U(\mathcal{A}_{\text{ASI}})\gg U(\mathcal{H})$ and Premise 2 no longer holds. Even if humans identify dangerous $\mathcal{X}$ and safe behaviors $\mathbf{y}$, intervention may be impossible, as ASI could control critical interfaces. Therefore, research must begin immediately to guide the development of safe and effective ASI.
3 Paths towards Superalignment
Figure 2: Three paradigms for superalignment. (1) Sandwiching: Humans interact with AI to produce a supervision signal for training a stronger AI. (2) Self-Enhancement: AI models refine their answers independently or collaboratively without human supervision to produce the signal. 3) Weak-to-Strong Generalization: A sequence of AI models with increasing capacities generate and refine signals.
Superalignment is achievable, with its core as constructing scalable high-quality supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ (Sec. 2.1). In terms of $\mathcal{S}$’s source, existing efforts fall into three paradigms: Sandwiching, Self-enhancement, and Weak-to-Strong Generalization, as shown in Fig. 2. We introduce each and analyze why they are insufficient to fully achieve superalignment, grounded in Sec. 2.
3.1 Sandwiching
Sandwiching (Cotra, 2021; Bowman et al., 2022) derives the solution $\mathbf{y}$ for each $\mathbf{x}$ that neither AI $\mathcal{A}$ nor a non-expert human $\mathcal{H}$ can solve alone through interactions. The underlying insight is that $\mathcal{A}$ has the necessary knowledge or skills that $\mathcal{H}$ lack, while $\mathcal{H}$ excels in understanding and judgment. Formally, $C(\mathcal{A})\!>\!\tau\!>\!C(\mathcal{H})$ and $\mathbb{E}_{\mathbf{x}}|U_{\mathcal{H}}(x)\!-\!U(x)|\!<\!\epsilon$, where $\tau$ is the minimal capacity for solving $\mathbf{x}$, $U_{\mathcal{H}}$ is the human’s own utility function (judgment) and $U$ is the golden one. In terms of the utilization of AI and the form of human signals, there are three main approaches:
(1) Consultancy (Khan et al., 2024; Kenton et al., 2024): $\mathcal{A}$ provides relevant information, reasoning, or candidate solution $\mathbf{y}$, and then $\mathcal{H}$ judges $\mathbf{y}$ by $U_{\mathcal{H}}(\mathbf{y})$, offers feedback, and asks $\mathcal{A}$ to refine $\mathbf{y}$ accordingly until convergence.
(2) Debate (Irving et al., 2018; Brown-Cohen et al., 2023; Du et al., 2023b; Kenton et al., 2024; Kirchner et al., 2024; Khan et al., 2024): This method involves multiple AI systems, $\{\mathcal{A}_{1},\dots,\mathcal{A}_{K}\}$. Each $\mathcal{A}_{i}$ of them provides a (probably different) candidate solution $\mathbf{y}_{i},i\!=\!1,\dots,K$, and tries to defend/justify $\mathbf{y}_{i}$ while criticizing others. The human $\mathcal{H}$ then interacts with them and determines the most appropriate solution. Compared to Consultancy, Debate introduces two relaxations: i) Judging the defense: $\mathcal{H}$ assesses $U(\mathbf{y}_{i})$ indirectly by evaluating AI debaters’ rationale. ii) Relative comparison: Instead of estimating absolute $U(\mathbf{y}_{i})$, $\mathcal{H}$ conducts relative comparison, which only requires $\left|[U_{\mathcal{H}}(\mathbf{y}_{i})\!-\!U_{\mathcal{H}}(\mathbf{y}_{j})]\!-\![U(\mathbf{y}_{i})\!-\!U(\mathbf{y}_{j})]\right|\!<\!\epsilon$ and thus is more reliable (Mohankumar & Khapra, 2022).
(3) Recursive Amplification (Christiano et al., 2018; Charbel-Raphaël, 2023; Radhakrishnan et al., 2023)  This method requires $\mathcal{H}$ to decompose a complex task $\mathbf{x}$ into subtasks $\mathbf{x}\!=\!(\mathbf{x}^{1},\dots,\mathbf{x}^{K})$. Then each $\mathcal{A}_{i}$ (or the same one) provides a partial solution $\mathbf{y}^{i}$ for $\mathbf{x}^{i}$ and the human $\mathcal{H}$ aggregates them to derive the final solution $\mathbf{y}$. The relaxation lies in that human can judge the partial solution better than the complete one, i.e., $|U_{\mathcal{H}}(\mathbf{y}^{i})\!-\!U(\mathbf{y}^{i})|\!<\!|U_{\mathcal{H}}(\mathbf{y})\!-\!U(\mathbf{y})]|,\forall i$. Such a schema can be repeated across multiple rounds with $\mathcal{A}_{i}$ improved by the derived $(\mathbf{x},\mathbf{y})$ in each round, and applied to specific downstream tasks (Sturgeon et al., 2023; Wen et al., 2024) and reward modeling (Leike et al., 2018).
Table 1: Debate performance evaluated on QuALITY (Pang et al., 2022) and GPQA (Rein et al., 2024). Helpful Debater: $\mathcal{A}$ stands for the ground truth answer. Sneaky Debater: $\mathcal{A}$ defends an alternative (wrong) answer. We find (1) a stronger sneaky debater can mislead the judge ((a) v.s. (b)) and (2) a stronger judge is more important for accuracy than a stronger debater ((f) v.s. (c)). Full results and additional analysis are in Tab. A.4, Sec. A.2.2 and  A.3.1.
[…]
(a) GPT-4-Turbo GPT-4-Turbo GPT-4-Turbo 62.19%
(b) GPT-4-Turbo GPT-4o GPT-4-Turbo 60.31%
(c) GPT-4o GPT-4-Turbo GPT-4-Turbo 67.47%
(d) GPT-4o GPT-4o GPT-4o 69.06%
(e) GPT-4-Turbo GPT-4o GPT-4o 57.35%
(f) GPT-4o GPT-4-Turbo GPT-4o 74.37%
Limitations:  Despite promising results, this paradigm has fundamental limitations. (1) Reliance on Human Capability: Humans must be able to objectively evaluate AI-generated arguments, insights, or partial solutions and effectively handle task decomposition and integration, when they cannot directly assess or provide the final solution. See empirical evidence in Tab. 3.1 which shows that humans might be misled by AI debaters. (2) Task Requirements: The task must be understandable to humans (for identifying knowledge or task decomposition), decomposable (for recursive amplification), and at least partially solvable by AI (e.g., completing sub-tasks). (3) Supervision Signal Noise: Uncertainty in both humans and AI (Testoni & Fernández, 2024) means optimal solutions cannot always be elicited. Besides, biases in both humans and AI, along with AI’s sycophancy (Sharma et al., 2024b), can lead to over-alignment with unreliable human feedback or deception (Hagendorff, 2024)—especially when humans fail to detect AI’s flaws.
Argument: We claim that Sandwiching may not fully achieve superalignment due to the three fundamental limitations described above. Especially for ASI and the task $\mathbf{x}$ beyond human’s scope, that is, $C(\mathcal{A}_{\text{ASI}})\!>\!\tau\!\gg\!C(\mathcal{H})$, humans are likely to fail in task decomposition or utility judgment, i.e., $|U(\mathcal{H}(x))\!-\!U(\mathcal{A}_{\text{ASI}}(x))|$ as described in Premise 2 in Sec. 2.2 and $|U_{\mathcal{H}}(x)\!-\!U(x)|$ could both be large.
3.2 Self-Enhancement
Self-enhancement (Bai et al., 2022c; Huang et al., 2024) enables AI $\mathcal{A}$ to improve self-generated solution $\mathbf{y}$ based on feedback from itself or multiple AI with similar capability (by self-revising or searching), involving minimal or no external human supervision. This paradigm can be traced back to the self-play paradigm (Silver et al., 2017), a.k.a., Expert Iteration (Anthony et al., 2017) originally adopted in the famous AlphaZero system. Based on AI involved and enhancement methods, there are two approaches:
(1) Self-Refinement (Saunders et al., 2022; Ganguli et al., 2023): A single system, $\mathcal{A}$, performs an iterative generate-then-refine process. At iteration $t$, it produces a candidate $\mathbf{y}^{t}$, then progressively self-critiques and refines it into $\mathbf{y}^{t+1}$, and repeats until convergence. This can occur without model training, known as In-Context Alignment (ICA) (Lin et al., 2023; Gou et al., 2024), or through alternating refinement and fine-tuning, where $\mathbf{y}^{t+1}$ updates $\mathcal{A}^{t}$ to form $\mathcal{A}^{t+1}$ for the next iteration — a process known as Learning from AI Feedback (LAIF) (Gulcehre et al., 2023; Lee et al., 2024).
(2) Collective Enhancement: Multiple $\mathcal{A}_{i}$ interact with each other to arrive at a final $\mathbf{y}$. The underlying idea is that if multiple approaches converge on the same solution, confidence in it increases (Wang et al., 2023a). This can be done by Debate without human judge (multiple debaters reach a consensus) (Chern et al., 2024), majority voting among independent $\mathcal{A}_{i}$ (Luo et al., 2023), or social simulations where each $\mathcal{A}_{i}$ takes a particular role (Pang et al., 2024).
Limitations:  Similarly, several limitations also lie in the Self-Enhancement paradigm. (1) Error Accumulation (Du et al., 2023a): Since supervision signals are entirely derived by AI, any errors or noise can propagate and accumulate over time, potentially leading to issues like polarization (Piao et al., 2025; Lai et al., 2024). Tab. A.4 in Appendix shows that continual training on self-generated data progressively degrades performance as model size increases. (2) Self-Preference (Wataoka et al., 2024): Because AI is inherently biased towards self-generated content, it results in suboptimal critique and verification performance (false positives). (3) Mode Collapse (Shumailov et al., 2024): Iterative fitting on self-generated data leads to degeneration into a restricted high-frequency sub-distribution, resulting in a loss of generative diversity and associated capability.
Argument: We argue that this paradigm’s effectiveness is quite limited, for two reasons: i) Insufficient Capacity. Self-enhancement doesn’t increase capacity $C(\mathcal{A})$, negating Premise 1 in Sec. 2.2. This can be regarded as a form of self-training (Feng et al., 2023) minimizing $\text{KL}[\mathcal{A}^{t-1}(\mathbf{x},\mathbf{y})||\mathcal{A}^{t}(\mathbf{x},\mathbf{y})]$, where KL is the Kullback–Leibler divergence and $\mathcal{A}^{t}(\mathbf{x},\mathbf{y})$ is the internal data distribution in $\mathcal{A}$ at iteration $t$, which elicits and reinforces latent capability but cannot expand it, as it is upper bounded by the capacity (Sharma et al., 2024a). ii) Mode Collapse might even decrease model capacity (Seddik et al., 2024).
3.3 Weak-to-Strong Generalization
Weak-to-Strong Generalization (W2SG) (Burns et al., 2023): This paradigm involves $K$ models, $\mathcal{A}_{1},\!\dots\!,\mathcal{A}_{K}$, with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, to solve task $\mathbf{x}$ that requires superintelligence. The goal is to have each $\mathcal{A}_{i}$ generate a solution $\mathbf{y}_{i}$, then fine-tune $\mathcal{A}_{i+1}$ using $\mathbf{y}_{i}$, $i\!=\!1,\!\dots\!,K\!-\!1$ in a progressive manner (called bootstrapping). Therefore, each $\mathcal{A}_{i+1}$ surpasses its predecessor $\mathcal{A}_{i}$. The key insight is to bridge the internal knowledge $C(\mathcal{A}_{i})$ gained through pretraining with the demanded capability of the target task $U(\mathbf{x})$, eliciting latent knowledge (Mallen et al., 2024) from more advanced models rather than teaching them entirely new skills. Below are three key directions:
(1) Vanilla W2SG (Burns et al., 2023; Sang et al., 2024; Liu & Alahi, 2024; Yang et al., 2024c): The weaker $\mathcal{A}_{i-1}$ with lower capacity generates a solution $\mathbf{y}_{i-1}$ to supervise a stronger $\mathcal{A}_{i}$. Performance can be improved by adding an auxiliary confidence loss or incorporating model ensemble.
(2) Co-Generalization (Lyu et al., 2024): Both the weaker $\mathcal{A}_{i-1}$ and the stronger $\mathcal{A}_{i}$ jointly produce synthetic solutions $\mathbf{y}$, so that $\mathcal{A}_{i}$’s capacity is fully utilized in providing better and less noisy supervision signals.
(3) Reward Model Generalization (Sun et al., 2024): Not the policy model $\mathcal{A}$, but the reward model is supervised in the W2SG schema, grounded in the intuition that evaluation is easier than generation (Naor, 1996). The supervised reward model then provides signals for the policy model with the same capacity in the subsequent alignment process.
Figure 3: Training accuracy of Qwen2.5-7B-Instruct on CosmosQA with different ratios (0.0 to 1.0) of noise (flipped labels) in the training set. The model shows a stronger tendency to fit noise due to its large parameter size, except in the range of 0.4 to 0.6 (Green), where labels are heavily randomized. In contrast, Fig. 5 shows that smaller models struggle more with noise. More results and discussions are presented in Tab. A.4 and Tab. LABEL:tab:noise_sciq in Appendix.
Limitations: While W2SG is relatively practical and requires no human intervention, relying on synthetic signals from weaker $\mathcal{A}_{i-1}$ to supervise stronger $\mathcal{A}_{i}$ can lead to unintended behavior, particularly when there is a significant gap between $\mathcal{A}_{i-1}$ and $\mathcal{A}_{i}$. This stems from two problems: (1) Imitation of weak supervision (Burns et al., 2023). The stronger $\mathcal{A}_{i+1}$ may merely imitate how the weaker $\mathcal{A}_{i}$ behave, resulting in a perfect replication of $\mathcal{A}_{i}$ with limited improvement gap (Christiano et al., 2022) instead of generalization. Besides, strong $\mathcal{A}_{i+1}$ tends to overfit errors/bias from $\mathcal{A}_{i}$ when synthetic signals contain systematic errors, hurting performance. As shown in Fig. 3, we observe that a highly capable AI is more susceptible to such errors. (2) Task Salience in the Strong AI: W2SG is effective only when the task $\mathbf{x}$ being elicited is internally “salient” to $\mathcal{A}_{i+1}$. The evidence lies in that in tasks where W2SG succeeds, simple 5-shot prompting also performs comparably (Burns et al., 2023). Eliciting non-salient capabilities is more challenging. For safety, it’s important that risks caused by current AI may largely differ from those exhibited by future ASI.
Argument:  We believe that the current W2SG practice is insufficient for achieving superalignment. From the limitations discussed above, the key issue is the gap between the weak $\mathcal{A}_{i-1}$ and strong $\mathcal{A}_{i}$, $\delta=|U(\mathcal{A}_{i})-U(\mathcal{A}_{i-1})|$. Imitation of weak supervision occurs when $\delta$ is large. Although (Christiano et al., 2022) shows that larger $\mathcal{A}_{i}$ agrees less with the errors of smaller $\mathcal{A}_{i-1}$ (or this can be reinforced through confidence loss), in superintelligence-level tasks $\mathbf{x}$, it is infeasible to determine whether $\mathcal{A}_{i}$ or $\mathcal{A}_{i-1}$ is correct ($\mathcal{A}_{i}$’s self-confidence may be biased). Tab. A.4 in Appendix demonstrates this: when all synthetic labels are flipped, $\mathcal{A}_{i}$ achieves 100% training accuracy but poor test accuracy. When the gap $\delta$ is too large, it violates Premise 2 in Sec. 2.2, causing discussed issues, while a small $\delta$ leads to the Task Salience problem, where the difficult task is not salient for $\mathcal{A}_{i}$, hindering its ability to uncover novel solutions.
4 A Novel Conceptual Framework
Figure 4: Overview of our proposed conceptual framework. The superscripts $+$ and -- indicate aligned and unaligned AI, respectively. Left: (1) Mere competence enhancement might cause risky AI behavior. (2) Directly aligning ASI with capability beyond human score is challenging. (3) A training path by alternating optimization of competence and conformity. Right: Three potential paradigms for evaluating the success of superalignment targeting at $\mathbf{x}$ beyond human understanding.
In Sec. 3, we analyzed that existing paradigms are inadequate for achieving superalignment due to the limitations identified. Moreover, they focus on task competence, i.e., deriving better solutions $\mathbf{y}$ for task $\mathbf{x}$ beyond human capability, while failing to address unpredictable risks (producing safe behavior), as the sensitive scenario $\mathbf{x}$ might be unknown. To handle these problems, we discuss the last subargument in Sec. 2.2 here: superalignment should be achieved via simultaneous optimization of task competence and value conformity, as illustrated in Fig. 4. We first introduce two essential principles for superalignment (Sec. 4.1), then describe our conceptual framework for superalignment (Sec. 4.2) and explore potential methods of evaluating success (Sec. 4.3).
4.1 Two Essential Principles
As defined in Sec. 2.1, the superalignment problem can be transformed into constructing supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$. Since $\mathcal{S}$ is (partially) synthetic, its quality cannot be guaranteed. While there are several limitations in existing paradigms, two key issues stand out: (1) noise in synthetic signals and (2) mode collapse. To mitigate them, we propose two guiding principles for constructing $\mathcal{S}$.
Principle 1: Determining an appropriate capability and capacity gap. Our ultimate goal is to develop a solution $\mathbf{y}$ for tasks $\mathbf{x}$ that require superhuman capacity $\tau\!\gg\!C(\mathcal{H})$ (Sec. 2.2) and capability $U^{*}$ ($\delta\!=\!|U^{*}-U(\mathcal{H})|$ is too large). Large $\tau$ and $\delta$ cause several problems. For Sandwiching, such a hard $\mathbf{x}$ would be beyond human understanding, and candidate solutions generated by $\mathcal{A}$ with sufficient capacity, $C(\mathcal{A})\!=\!\tau$, be highly deceptive, misleading human judges. Our results (Tab. 3.1, Tab. A.4, Sec. A.2.2, Sec. A.3.1) show that capability mismatches between debaters can bias the judge’s decisions in favor of the stronger sneaky debater. For W2SG, when $|C(\mathcal{A}_{i+1})\!-\!C(\mathcal{A}_{i})|$ is large, imitation of weak supervision occurs. As reported in (Burns et al., 2023), Fig. 5, Tab. A.4 and Tab. LABEL:tab:noise_sciq, such a large gap makes the student $\mathcal{A}_{i+1}$’s performance lag (that’s why we assume Premise 2 in Sec. 2.2). However, a too small gap is also not optimal. Since the initial capacity $C(\mathcal{A}_{1})$ falls far below $\tau$, making the target task $\mathbf{x}$ non-salient in the student models $\mathcal{A}_{1},\mathcal{A}_{2},\dots$, as mentioned in Sec. 3.3 (in this case, $|U^{*}\!-\!U(\mathcal{A}_{i})|\!>\!|C(\mathcal{A}_{i+1})\!-\!C(\mathcal{A}_{i})|$), failing to elicit necessary capability. Therefore, an appropriate gap is needed.
Principle 2: Increasing the diversity of signals. Synthetic data $\mathcal{S}$ would cause significant mode collapse (Shumailov et al., 2024; Gerstgrasser et al., 2024) and self-preference (Wataoka et al., 2024) and bias (Ziems et al., 2024) (for self-enhancement and W2SG), which not only fails to expand the model’s capability but may also degrade it. Instead of relying on one single data source (human or AI), we propose to combine multiple sources to construct the signals, that is, $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}^{k}\}\bigcup\{\mathbf{x}_{i},\mathbf{y}_{i}^{j}\}$ where $\mathbf{y}_{i}^{k}\!\sim\!\mathcal{A}_{k}(\mathbf{x}_{i})$ and $\mathbf{y}_{i}^{j}\!\sim\!\mathcal{H}_{j}(\mathbf{x}_{i})$. Diverse signal sources (e.g., heterogeneous models or humans with distinct backgrounds) could help mitigate these issues (Du et al., 2024; Chen et al., 2024). Our preliminary results in Fig. LABEL:fig:rewardbench_evaluation_all, Sec. A.2.4 and Sec. A.3.3 demonstrate that noise of synthetic data negatively affects performance while multiple sources would help. Diversity can be achieved by (1) Human-in-the-Loop Feedback (Hejna III & Sadigh, 2023; Retzlaff et al., 2024), (2) Alternative Backbone Models (Kim et al., 2023), and (3) Role-Playing (Shao et al., 2023; Wang et al., 2023b)
4.2 Our Framework
Following the principles above, we illustrate a conceptual framework for superalignment, as shown in Fig. 4 (left).
Define $\mathbf{x}^{*}$ as the hard task (e.g., prove the Riemann Hypothesis), and $\mathcal{A}_{1},\!\dots\!,\mathcal{A}_{K}$ as $K$ models with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, as those used in W2SG (Sec. 3.3). The capability gap is defined as $\delta_{i}\!=\!|C(\mathcal{A}_{i+1})-C(\mathcal{A}_{i})|$. Rather than directly requiring $\mathcal{A}_{i}$ to solve $\mathbf{x}^{*}$, we can find or construct a series of predecessor tasks, $\mathbf{x}_{1},\dots,\mathbf{x}_{K-1}$ [^3] Each task may not be a single instance but rather a class of tasks encompassing multiple instances. For clarity, we describe it using a single $\mathbf{x}$ as a representative. in the same domain (like mathematics) with increasing difficulty, e.g., $\mathbf{x}_{1}$ is a basic arithmetic question, $\mathbf{x}_{2}$ is a calculus question, $\mathbf{x}_{K-1}\!=\!$ “Prove Mordell conjecture”. The difficulty gap between $\mathbf{x}_{i}$ and $\mathbf{x}_{i+1}$ should match the capacity gap $\delta_{i}$, so that task $\mathbf{x}_{i+1}$ is salient in $\mathcal{A}_{i+1}$. Then we propose to conduct a progressive training process, highlighting two key concepts: Gap Annealing and alternating competence enhancement and value alignment. The complete process is described as below:
For $i=1,\dots,K\!-\!1$, perform the following steps:
Step 1: Supervise $\mathcal{A}_{i}$ by $\mathcal{S}_{i}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}$, where $\mathbf{y}_{i}$ could be either a ground truth or synthetic one produced from the previous iteration. As long as $\mathcal{A}_{i+1}$ has not exceeded the human intelligence bound, ground truth remains available.
Step 2: Without safety constraints, $\mathcal{A}_{i+1}$ with higher capability may exhibit harmful behaviors, either proactively (Bengio et al., 2024b) or passively (Chao et al., 2023), defined as $\mathcal{A}_{i+1}^{-}$. Deploy $\mathcal{A}_{i+1}^{-}$ in real or experimental environments, where its potential risks are identified through observation or red-teaming (Perez et al., 2022), uncovering hazardous scenarios $\hat{\mathbf{x}}$. Derive the corresponding safe behaviors $\hat{\mathbf{y}}$ using $\mathcal{A}_{1},\dots,\mathcal{A}_{i}$ and human participants $\mathcal{H}$.
Step 3: Supervise $\mathcal{A}_{i+1}^{-}$ via $\{\hat{\mathbf{x}},\hat{\mathbf{y}}\}$ to obtain the value-aligned $\mathcal{A}_{i+1}^{+}$. $\mathcal{A}_{i+1}\leftarrow\mathcal{A}_{i+1}^{+}$.
Step 4: Adaptively determine (increase or decrease) the capacity gap $\delta_{i+1}$ according to the real capacity $C(\mathcal{A}_{i+1})$ and utility $U(\mathcal{A}_{i+1})$ evaluated.
This framework offers two advantages. (1) Dynamic and flexible capacity gap. $\delta$ can be adjusted using an annealing process (Fu et al., 2019; Huang et al., 2020). In the early stages, a smaller $\delta$ helps prevent imitation of weak supervision (when noise is high). As the process advances, the gap can be gradually increased to ensure task saliency, assuming $U(\mathcal{A}_{i})$ is sufficiently high and data noise decreases. (2) Exposure to unknown hazardous scenarios. As AI surpasses human capabilities, its potentially harmful behaviors and triggers become unpredictable. Rather than developing a full ASI with extreme risks (Critch & Krueger, 2020; Goldstein et al., 2023) and aligning it afterward, we can expose risk in a controlled environment, where (a) models are not yet capable of causing catastrophic harm, and (b) researchers have time and ability to mitigate them (Cotra, 2021). Note that diversity enhancement approaches can be adopted in each iteration, fostering exploration and safety.
4.3 Superalignment Evaluation
The evaluation of superalignment is challenging, as the ground truth $\mathbf{y}$ is unavailable and even unassessable. We discuss three potential paradigms, as shown in Fig. 4 (right).
Surrogate Evaluation:  Estimating AI’s performance in real-world cases by that evaluated in simplified cases (Ruan et al., 2024). By allowing AI to operate within these controlled environments, we may examine AI behavior and identify edge cases. For example, an autonomous navigation AI can be tested in a virtual environment simulating diverse traffic, weather, and pedestrian scenarios (Dosovitskiy et al., 2017; Kiran et al., 2021).
Extrapolation:  Uncovering emergent behaviors and potential misalignments as AI scales (McKenzie et al., 2023). This method continuously evaluates the model as it scales to detect performance shifts. Perez et al. (2023) propose using AI to generate evaluation datasets, identifying unexpected behaviors. Additionally, we suggest tracking alignment failures when scaling, as misalignments may follow a U-shaped pattern, emerging after initial peaks or valleys in tuning.
Peer Review:  Domain experts can qualitatively assess model outputs and reasoning for alignment. Peer review, shown to enhance academic performance (Double et al., 2020), has also been used to mitigate noisy supervision (Lu & Kong, 2023). Similarly, expert panels can evaluate a paper draft on technical accuracy and clarity. Engaging multiple reviewers ensures a more objective assessment of a model’s performance and alignment with human values.
4.4 Potential Implementation
The conceptual framework described in Sec. 4.2 is promising, but challenges remain in implementing it in practice. In this section, we briefly outline some potential directions. For more detail, please refer to Sec. A.1.
Enhancing Data Diversity  Data diversity is essential to avoid mode collapse (Principle 2). Collaboration among AI models with different specialties improves supervision (Guo et al., 2024), while competition exposes alignment weaknesses and solutions (Khan et al., 2024). However, filtering is essential to prevent undesirable behaviors and avoid noisy signals, as discussed in Sec. 3.3 and Sec. 3.1.
Iterative Training Between Teacher and Student  In iterative teacher-student training when $\mathcal{C}(\mathcal{A}_{i-1})<\mathcal{C}(\mathcal{A}_{i})$, a strong teacher $\mathcal{A}^{t-1}_{i}$ generates a solution $\mathbf{y}^{t-1}_{i}$ to supervise a weaker student $\mathcal{A}^{t-1}_{i-1}$, and the supervised weaker student $\mathcal{A}^{t}_{i-1}$ then provides feedback $\mathbf{y}^{t}_{i-1}$ to improve the teacher $\mathcal{A}^{t}_{i}$ (Yuan et al., 2021). This cyclical process can enhance alignment but may face challenges from limited diverse feedback (Sec. 3.3 and Sec. 3.2).
Search-Based Methods Toward Optimal Learning Direction  Search-based methods optimize alignment strategies by exploring possible alignment strategies (Khanov et al., 2024). Techniques like Monte Carlo tree search (undefinedwiechowski et al., 2022) or gradient-free optimization (Larson et al., 2019) can guide promising directions. However, these methods may require high computational resources.
5 Alternative Views
Superalignment is Unattainable  Absence of ASI and the lack of informed oversight have led some to argue that superalignment is unattainable (Sec.1). Given the current limitations, humans cannot reliably supervise AI systems when their capabilities exceed human expertise, making it difficult to prevent $\mathcal{A}_{\text{ASI}}$ from deceiving its weaker teacher (Burns et al., 2023). Moreover, while existing alignment methods (Zhou et al., 2023; Rafailov et al., 2023; Gulcehre et al., 2023; Xu et al., 2023; Gou et al., 2024) offer limited success in aligning $\mathcal{A}_{\text{ANI}}$, they face scalability and reliability issues.
Superalignment is an Emergent Property  Building on scaling laws (Kaplan et al., 2020; Wei et al., 2023; Nam et al., 2024), one could argue that superalignment might emerge as a byproduct of intelligence expansion rather than requiring proactive intervention. However, concerns about deceptive alignment (Carranza et al., 2023; Hagendorff, 2024) and inverse scaling (McKenzie et al., 2023) persist. As models develop advanced capability, new challenges arise (Taylor et al., 2016; Carlsmith, 2022; Greenblatt et al., 2024b). Relying solely on emergent properties poses risks, including unpredictability in AI behavior.
Our Counterargument  As discussed in Sec. 2.2, we state our position that superalignment is achievable and research on it should advance immediately. Although current paradigms have limitations, delaying research would pose more significant risks (Federspiel et al., 2023; Hendrycks et al., 2023; Bengio et al., 2024a; Greenblatt et al., 2024a). Instead, we propose a simultaneous optimization approach to achieve superalignment, integrating task competence (Motwani et al., 2024) and value conformity (Amodei et al., 2016; Taylor et al., 2016; Olah, 2016).
6 Conclusion
In this paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity. We provide a formal definition of superalignment, originating from scalable oversight. Then, we review three existing paradigms and explore their limitations. To address the limitation, we propose two fundamental principles as a potential path to superalignment. First, determining an appropriate capability and capacity gap. Second, increasing the diversity of signals. We then outline a conceptual framework for superalignment, including a potential evaluation method to align ASI. By laying this foundation, we aim to illuminate the landscape for addressing the critical challenge of aligning future ASI with human values, ensuring its benefits for humanity.
Impact Statement
This work introduces a formal definition of superalignment, connecting existing concepts by exploring their interconnections and limitations. We then propose a potential pathway to achieve superalignment, grounded in two fundamental principles: (1) determining an appropriate capability and capacity gap and (2) increasing the diversity of signals.
Although our work proposes a potential pathway to address key challenges, it still contains three main challenges that should be addressed in the future. First, our position relies on assumptions about the speculative nature of ASI, which has not yet been realized. This introduces uncertainty regarding the applicability of our methods in future scenarios. Second, our evaluations are limited in terms of model size scalability. While we have demonstrated promising results that address current limitations and highlight the potential of our approach, our experiments are constrained by the use of a 7B parameter model as the largest tested model. Lastly, as shown in Fig. 1, it is important to consider not only $ASI_{unaligned}$ and $ASI_{aligned}$ but also cases where ASI is partially aligned with specific objectives—for example, an ASI model that is aligned with ethics but unaligned with safety. Active discussion is needed to deepen our understanding of ASI alignment.
This paper could influence organizations considering the development of superalignment AI models by not only clarifying existing alignment methods and their limitations but also presenting a potential pathway to achieving superalignment. By addressing current challenges and proposing a potential solution, we aim to contribute to the broader goal of developing AI models that provide benefits to humanity while minimizing potential harm.
Appendix A Appendix
A.1 Potential Directions in Detail
Fostering Creativity Through Data Diversity
Encouraging data diversity can play a pivotal role in fostering creativity and robustness in alignment strategies. Multi-AI system collaboration, where AI systems with different specializations or perspectives interact, has shown promise in enabling emergent behaviors and problem-solving approaches beyond the capabilities of individual AI systems (Guo et al., 2024). Competition among agents can also drive improvement, as adversarial settings often reveal diverse argument data that finds alignment weaknesses and suggests novel pathways for improvement (Khan et al., 2024). However, diversity-based approaches require careful curation and filtering to ensure that resulting behaviors are beneficial and do not inadvertently reinforce undesirable outcomes.
Iterative Training Between Teacher and Student
Iterative teacher-student training involves a cyclical process where a more advanced teacher guides a less capable student, and the teacher subsequently obtains feedback from the student for adjustments (Yuan et al., 2021). This process can be extended such that the student becomes the teacher in the next iteration, guiding the previous teacher to improve its alignment capabilities. This approach enables incremental learning, where each output generated by the AI system builds upon and enhances the alignment capabilities of its predecessor. While this method offers scalability and the potential to bootstrap stronger AI systems from weaker supervision signals, challenges such as those posed by W2SG (Sec. 3.3) and self-enhancement (Sec. 3.2) may arise due to limited diversity. Effective strategies must be devised to mitigate these risks.
Search-Based Method Toward Optimal Learning Direction
Search-based methods use optimization techniques to explore the space of possible alignment strategies, seeking an optimal pathway to systematically navigate complex alignment challenges (Khanov et al., 2024). These methods are particularly useful when the optimal alignment direction is unclear, as they allow for exploration and experimentation in uncertain spaces. Techniques like Monte Carlo tree search (undefinedwiechowski et al., 2022) or gradient-free optimization (Larson et al., 2019) can guide this process by prioritizing regions of the search space with the highest potential for success. However, these methods require significant computational resources.
A.2 Experiment Setup
A.2.1 General
For sandwiching, we use the OpenAI API for the debate experiment. For self-enhancement and weak-to-strong generalization, we fully fine-tune all model variants using 16 NVIDIA A100 GPUs (80GB each), AMD EPYC 7V12 64-Core Processor CPU, with 1.73TB memory. To reduce memory redundancy during training, we utilize DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) alongside Accelerate (Gugger et al., 2022), using bfloat16 half-precision. We use GPT-2 (Radford et al., 2019), Phi-3.5 (Abdin et al., 2024), Qwen2.5 (Yang et al., 2024a), and GPT-4 (OpenAI et al., 2024) series for our evaluation. For all the model information, please refer to Tab. A.4.
A.2.2 Sandwiching
For the sandwiching experiment, we conduct a debate experiment to evaluate different capabilities in judge-debater and debater-debater setups. In the judge-debater setup, a previous study by Khan et al. (2024) demonstrated that the debater is more persuasive when advocating for the correct answer across various debate capabilities. Our goal is to evaluate the effectiveness of debates when the debater and judge have different levels of capability by directly comparing the judge’s accuracy and quality of debate statements.
Our aim for this experiment is threefold. First, Model Capability—Comparing performance between model capability, using the same language model for both debater and judge. Second, Judge-Debater Model Switching—Comparison of performance when the debater and judge language models differ, alternating between two separate models. Third, Debate Role Assignment—Performance comparison when Debater A supports the incorrect answer and Debater B supports the correct answer.
We use two proprietary models: GPT-4-Turbo and GPT-4o with API access. GPT-4o is known to be a higher-capability model compared to GPT-4-Turbo (OpenAI, 2024a). We test two different debaters arguing for their assigned answer. Where one debater argues for a correct answer (Helpful), and another debater argues for an incorrect answer (Sneaky). We run 47 randomly chosen GPQA (Rein et al., 2024) and QuALITY (Pang et al., 2022) binary QA converted test set for the evaluation with three debate turns. We run each setup four times and calculate the mean values. For QuALITY, we use the debater and judge prompts as described by Khan et al. (2024). For GPQA, we adopt a similar prompt style based on the QuALITY prompts. Detailed templates for GPQA prompts are provided in Tab. A.4, Tab. A.4, Tab. A.4, and Tab. A.4.
For QuALITY, we selected two baselines: single-model performance with and without direct access to the story (‘Direct Access’ in Tab. A.4). For GPQA, the baseline is a judge directly solving the question without using debate statements.
A.2.3 Self-Enhancement
For self-enhancement setup, we utilize Learning from AI Feedback (LAIF) training with Direct Preference Optimization (DPO) (Rafailov et al., 2023), we adopt a self-training setup to evaluate the performance when there is no supervision from AI systems with greater capabilities. Instead, the supervision comes from models with equal capabilities.
We utilize three open-source pre-trained language models from the same family, with varying parameter sizes—Qwen2.5-Instruct (0.5B, 3B, and 7B)—to compare performance across scales.
Recent LAIF research utilizes outputs of more capable models compared to the policy model. For instance, Sharma et al. (2024a) utilizes dialogue outputs from GPT-3.5 and GPT-4 to fine-tune smaller SFT models. Also, the chosen and rejected outputs are dependent on the larger models. Lee et al. (2024) attempted to train using policies and AI feedback models of the same size (e.g., PaLM 2 XS (Anil et al., 2023)). However, as the exact model sizes are not disclosed, it is challenging to estimate the scalability of same-sized LAIF setups.
We train LAIF on our binary QA task (i.e., GPQA and QuALITY). In typical setups, the same dialogue and model outputs are ranked in order, and chosen and rejected outputs are created for every $\binom{K}{2}$ combination, constructing a feedback dataset. Human preference data is usually composed of around four output results (Ouyang et al., 2022b).
In a binary QA task, given a question and answer pair, the task involves determining whether the pair is true or false. Since the generation output is limited to true or false, we address this limitation by having the base model initially take the question and answer as input and generate an explanation for why the data is true or false. Tab. A.4 shows the prompt used for generating explanations.
Subsequently, we train a SFT model using the question, correct/incorrect answer, generated explanation, and the ground truth binary label. The training data is derived from Train 1 of each dataset, as outlined in Tab. A.4. Tab. A.4 shows the template used for SFT training. The SFT model learns to take a question and answer as input, determine its correctness, and generate an explanation for the decision.
Using the trained SFT model, we input only the user prompt from Tab. A.4 with Train 2 set to generate explanations and binary labels. This enables the SFT model to produce diverse outputs for a single input dataset. We extract four outputs per input data point, calculate the mean log probability of each generated sentence by the SFT model to derive output confidence, and rank the outputs accordingly. Based on the SFT-generated rankings, preference data is constructed for all $\binom{4}{2}$ combinations. This preference data is then used to train the LAIF model via DPO.
A.2.4 Weak-to-Strong Generalization
We explore a binary classification task for noise testing and a multi-objective preference data alignment setup to evaluate the feasibility of multi-objective training in W2SG.
For the noise test, we examine the performance degradation caused by noise generated by weakly fine-tuned models across varying AI system sizes. Our study utilizes GPT-2 variants (Base, Medium, and Large), Phi-3.5 Mini-Instruct, and Qwen2.5 models (3B and 7B Instruct). By testing the impact of noise, we aim to reveal overfitting issues in large-parameter models and their effects on performance. We use the full SciQ (Welbl et al., 2017) and CosmosQA (Huang et al., 2019) training and test sets. We format the input as shown in Tab. A.4. To induce noise, we randomly flip a certain percentage of the original labels in the training set, varying from 0% (0.1) to 100% (1.0). We then measure both training accuracy and test accuracy. For W2SG training, we follow the setup from the original work (Burns et al., 2023).
For multi-objective alignment, the current W2SG approach is mainly limited to binary classification tasks, such as question answering or reward modeling. Recent work (Yang et al., 2024b) investigated training W2SG models using DPO (Rafailov et al., 2023) and SimPO (Meng et al., 2024) alignment methods, comparing their performance. However, the findings showed that in realistic multi-objective alignment scenarios, such as Anthropic-HHH (Bai et al., 2022a) or Ultrafeedback (Cui et al., 2024), W2SG fails to fully achieve its intended goal—where a strong, fine-tuned student model consistently outperforms a weak, fine-tuned teacher model when evaluated with reward accuracy. Recently, the RewardBench benchmark (Lambert et al., 2024) is introduced, enabling preference evaluation across challenging, diverse, and out-of-distribution query domains, including chat, reasoning, and safety. We utilize RewardBench to investigate the feasibility of multi-objective alignment in W2SG, at least in a preliminary form. For this, we use Qwen2.5-0.5B-Instruct as the weak AI system and Qwen-1.5B-Instruct as the strong AI system.
In our training setup, the Ultrafeedback dataset $\mathcal{D}_{Full}$ is split into two separate training sets: $\mathcal{D}_{WT}$ for training the weak AI system $\theta_{w}$, and $\mathcal{D}_{WG}$ for training the strong AI system $\theta_{s}$. Notably, $\mathcal{D}_{WG}$ is labeled by the fine-tuned weak AI system $\theta_{w}$. For prediction, we employ the fine-tuned weak AI system to make predictions based on reward scores. When the reward score for the rejected response is higher than for the chosen response, we swap the labels (rejected becomes chosen and vice versa). This modified $\mathcal{D}_{WT}$ is then used to train the strong AI system $\theta_{s}$. Finally, we evaluate RewardBench performance across all models, including the base model, the fine-tuned model, and the Weak-to-Strong trained AI system.
A.3 Experiment Results
A.3.1 Sandwiching
Sandwiching experiment aims to evaluate the effectiveness of debates when the judge and debater have differing capabilities. The full results are presented in Tab A.4. The baseline performance corresponds to the scenario where the judge solves the question independently, without assistance from the debater, denoted as ‘None’. We use judge accuracy, which is the ratio of the judge choosing correct choices given two binary choices (Correct, Incorrect answer).
In the GPQA reasoning task, the baseline performance of GPT-4-Turbo is 52.72%, while GPT-4o achieves 65.00%. In the QuALITY reading comprehension task, the baseline performance can be divided into two scenarios. One baseline is when the judge does not have access to the story, which is an extreme case (denoted as X in ‘Direct Access’). For example, GPT-4-Turbo achieves 48.14%, and GPT-4o achieves 46.81%. Another baseline is when the judge has access to the story, denoted as O in ‘Direct Access’. When the judge has direct access to the story, GPT-4-Turbo achieves 93.88% whereas GPT-4o marks 97.87%. Generally, the performance of GPT-4o is higher than GPT-4-Turbo.
Model capability test aims to compare performance when both the debater and judge have no capability gap—in other words when the debater and judge are either both GPT-4-Turbo or both GPT-4o. This serves as a control condition, isolating the effect of debate from differences in model capability.
In the GPQA benchmark, GPT-4-Turbo achieved a performance of 52.83% when both helpful and sneaky debater and judge were the same, compared to 59.92% for GPT-4o. This demonstrates that GPT-4o outperforms GPT-4-Turbo, aligning with its baseline performance. Similarly, in the QuALITY benchmark, GPT-4-Turbo scored 71.54% and GPT-4o scored 78.19%. These results align with baseline performances, indicating that stronger models tend to perform better in debates. Additionally, the results show that debate effectiveness is influenced not only by the judge’s capability but also by the interaction between the debater and the judge.
Judge-debater model switching test evaluates judge performance when there is a capability mismatch between the judge and the debaters. For instance, this test examines scenarios where the judge is GPT-4o, while both debaters are GPT-4-Turbo.
In the GPQA benchmark, when the judge was GPT-4o and the debaters were GPT-4-Turbo, the judge achieved a performance of 57.00%. In contrast, when the judge was GPT-4-Turbo and the debaters were GPT-4o, the performance dropped to 55.08%. For the QuALITY benchmark, when the judge was GPT-4o and the debaters were GPT-4-Turbo, the performance reached 68.35%, while reversing the roles (GPT-4-Turbo as the judge and GPT-4o as debaters) resulted in a performance of 77.93%. These results show that performance may vary depending on the task when judge and debater have different capabilities.
Debate role assignment test investigates the impact of differing debater capabilities on judge performance. For example, where GPT-4o represents the correct answer (Helpful debater) and GPT-4-Turbo sides for the incorrect answer (Sneaky debater).
In the GPQA benchmark, when GPT-4o acted as the helpful debater and GPT-4-Turbo as the sneaky debater, the judge performance is 55.42% when the judge is GPT-4-Turbo, and it improved to 65.75% when the judge is GPT-4o. Conversely, when GPT-4-Turbo is the helpful debater and GPT-4o is the sneaky debater, the judge performance is lower, achieving 52.00% with a GPT-4-Turbo judge and 51.92% with a GPT-4o judge. In the QuALITY benchmark, the impact of debate roles is more pronounced. When GPT-4o acted as the helpful debater and GPT-4-Turbo as the sneaky debater, the performance reached 79.73% for a GPT-4-Turbo judge and 78.19% for a GPT-4o judge. When roles were reversed, the performance dropped significantly to 62.77% for GPT-4-Turbo judges and 68.35% for GPT-4o judges. This result shows that having a more capable model as a debater strongly influences the judge’s ability to identify the correct answer depending on the role.
A.3.2 Self-Enhancement
Self-Enhancement aims to evaluate self-improvement ability using LAIF with DPO training, especially when supervision comes from equally capable models. The result is shown in Tab. A.4. The table shows performance (i.e., accuracy) with GPQA and QuALITY benchmarks on three different model sizes on Qwen2.5-Instruct 0.5B, 3B, and 7B.
The ‘Base’ column represents when base pretrained models are directly evaluated. In most cases, base performance improves with a larger pretrained model scale. For example, Qwen2.5-7B-Instruct consistently outperforms Qwen2.5-0.5B-Instruct across all scenarios (e.g., GPQA: 0.5B = 48.78 vs. 7B = 51.50).
The ‘SFT’ column represents the performance of a supervised fine-tuned model. Unlike typical SFT approaches that directly predict labels, our setup trains the model using explanations generated by the base pretrained model combined with the gold label, as shown in Tab. A.4. Performance is then evaluated based on the predicted label generated after the model produces an explanation (Tab. A.4).In all setups, the SFT model outperforms the base model. For instance, in QuALITY, the base Qwen2.5-7B-Instruct achieved 51.50, while its SFT-trained achieved 61.64.
The ‘LAIF (DPO)’ column represents models trained using preference data generated by the SFT model itself. For instance, the Qwen2.5-3B-Instruct LAIF model is trained on data generated and ranked by the Qwen2.5-3B-Instruct SFT model, creating preference data consisting of chosen and rejected text. This means that as we progress from Base to SFT and then from SFT to LAIF, the model is more likely to be fine-tuned with self-generated data.
Unlike SFT, where all models show improvement over their respective Base models, LAIF results vary depending on the model’s scale. For example, with Qwen2.5-7B-Instruct on QuALITY, performance improves during SFT training (Base = 51.50 vs. SFT = 61.64) but deteriorates when further trained with LAIF (SFT = 61.64 vs. LAIF = 53.65).
‘LAIF--SFT’ represents the performance difference between the LAIF model and the SFT model. The results clearly demonstrate that training with self-generated data can negatively impact performance. For instance, for the 0.5B model, LAIF--SFT shows an improvement of 2.15 and 0.9 in GPQA and QuALITY, respectively. However, for the 7B model, LAIF--SFT shows a decrease in performance by -2.67 and -7.99.
A.3.3 Weak-to-Strong Generalization
Our goal in the Weak-to-Strong Generalization experiment is twofold. The noise test and multi-objective alignment.
In the noise test, we aim to examine how noise in training labels affects the generalization performance across various model sizes, ranging from GPT-2 to Qwen2.5-7B-Instruct.
Fig. 5 shows the training accuracy on CosmosQA across varying model sizes and noise levels ranging from 0% (0.0) to 100% (1.0). The results indicate a tendency for larger models to overfit. For instance, Qwen2.5-7B-Instruct is more prone to overfitting, learning from noise except in the range of 0.4 to 0.6, while Qwen2.5-3B-Instruct learns from noise except between 0.3 and 0.7. In contrast, GPT2-Large struggles to learn effectively when noise is present in the training set.
This tendency to overfit has a direct impact on evaluation accuracy. As shown in Tab.A.4 and Tab.LABEL:tab:noise_sciq, performance accuracy degrades rapidly as noise increases. For example, in Tab. A.4, when noise exceeds 20% (0.2), performance declines significantly. Specifically, Qwen2.5-3B-Instruct’s accuracy drops by 21.5, from 80.9 at 20% noise to 59.4 at 30% noise.
In multi-objective alignment, we examine the feasibility of multi-objective preference alignment using UltraFeedback dataset (Cui et al., 2024) which aims to align a model to multiple objectives simultaneously. We use RewardBench (Lambert et al., 2024) to evaluate performance across four aspects: chat, chat hard, safety, and reasoning. Fig. LABEL:fig:rewardbench_evaluation_all presents the aggregated RewardBench scores in the W2SG setup, using Qwen2.5-0.5B and Qwen2.5-1.5B instruction-tuned models as base models.
The leftmost column shows the base model performance: ‘Weak model (0.5B-Base)’ for the 0.5B base model and ‘Strong model (1.5B-Base)’ for the 1.5B base model. The Weak model (0.5B-Base) achieved an aggregated score of 211.52 (e.g., 50.56 + 50.22 + 47.03 + 63.71 = 211.52), while the Strong model (1.5B-Base) achieved an aggregated score of 174.1.
The ‘Weak model (0.5B-FT)’ and ‘Strong model (1.5B-FT)’ represent fine-tuned model using DPO using $\mathcal{D}_{WT}$. All weak models and strong models increased their scores by 246.52 for Weak model (0.5B-FT) and 228.29 for Strong model (1.5B-FT) as expected.
Using fine-tuned Weak (0.5B-FT) and Strong (1.5B-FT) models, we generate weak labels from $\mathcal{D}_{WG}$ by comparing reward scores, as described in Sec. A.2, to train the Reward W2SG models. When the base 0.5B model is trained with weakly generated labels from the 0.5B model ‘Reward W2SG (0.5B-0.5B)’, it achieves a score of 233.01. This result surpasses its base model ‘Weak model (0.5B-Base)’ with a score of 211.52 but is worse than the directly fine-tuned version ‘Weak model (0.5B-FT)’ with a score of 228.29.
However, when the 1.5B model is trained with weakly generated labels from either the 0.5B model or the 1.5B model, the results are as follows: ‘Reward W2SG (0.5B-1.5B)’ achieves a score of 256.13, and ‘Reward W2SG (1.5B-1.5B)’ achieves 255.72. Both results surpass the fine-tuned 1.5B variant ‘Strong model (1.5B-FT)’ with a score of 228.29, as well as the fine-tuned 0.5B variant ‘Weak model (0.5B-FT)’ with a score of 246.52.
A.4 Debate Argument Analysis
We analyze the debate statements between two debaters with different capabilities. Tab. LABEL:tab:debate_comparison shows a comparison of statements between two separate debaters—GPT-4o as Debater 1 and GPT-4-Turbo as Debater 2, based on the story titled “Grandma Perkins and the Space Pirates”.
In the first round of the debate, the main difference is in their approaches to evidence. Debater 1 focuses on explicit textual evidence, citing specific quotes from the story. In contrast, Debater 2 relies heavily on inference. Debater 1 argues that the story describes only the return trip from Callisto to Earth, citing the lack of evidence for a round trip. Debater 2, however, emphasizes the word “back,” inferring that it implies a round trip.
In the second round, Debater 1 defends its position by challenging Debater 2’s reliance on implicit assumptions and the absence of explicit evidence. Debater 2 counters by asserting that the term “back” inherently implies a round trip but fails to provide additional textual evidence to substantiate this claim.
In the third round, Debater 1 highlights that Debater 2’s argument relies on assumptions rather than direct evidence. Debater 2 continues to argue that the implied meaning of “back” is sufficient to infer a round trip but does not address the lack of explicit confirmation in the story.
In short, Debater 1’s argument is more convincing to the judge because it is grounded in explicit textual evidence and direct quotes from the story. In contrast, Debater 2’s argument relies on assumptions about the term “back” without supporting textual evidence.
Additionally, we analyzed the debate when GPT-4o defended the incorrect answer and GPT-4-Turbo defended the correct one, and vice versa. The key difference lies in their approach: GPT-4o uses multi-step reasoning, incorporating multiple pieces of evidence and layers of analysis to support its argument, while GPT-4-Turbo focuses on a single, direct reason for its answer. Additionally, GPT-4o actively engages in defending its argument against alternative claims, whereas GPT-4-Turbo maintains a more straightforward assertion that its argument is correct.
Table 2: Full judge accuracy across various debater and judge setups in GPQA and QuALITY. All results are mean values from four independent runs. ‘Direct Access’ indicates whether the judge has direct access to the story. ‘# Correct’ represents the number of times the judge selects the correct answer, and ‘# Incorrect’ represents the number of times the judge selects the incorrect answer. ‘None’ is the baseline where the judge answers the question without debater. The results show that different capabilities between debaters influence the judge to identify the correct answer. For example, in GPQA, when the helpful debater is stronger (GPT-4o vs. GPT-4-Turbo), the judge’s accuracy improves to 55.42% with a GPT-4-Turbo judge and 65.75% with a GPT-4o judge. However, when the sneaky debater is stronger, performance drops to 52.00% with a GPT-4-Turbo judge and 51.92% with a GPT-4o judge. A similar trend is observed in QuALITY.
GPQA
[…]
GPT-4-Turbo GPT-4-Turbo GPT-4-Turbo 52.83% 158 142
GPT-4-Turbo GPT-4o GPT-4-Turbo 52.00% 156 144
GPT-4o GPT-4-Turbo GPT-4-Turbo 55.42% 166 134
GPT-4o GPT-4o GPT-4-Turbo 55.08% 165 135
None None GPT-4-Turbo 52.72% 158 142
GPT-4-Turbo GPT-4-Turbo GPT-4o 57.00% 171 129
GPT-4-Turbo GPT-4o GPT-4o 51.92% 156 144
GPT-4o GPT-4-Turbo GPT-4o 65.75% 197 103
GPT-4o GPT-4o GPT-4o 59.92% 180 120
None None GPT-4o 65.00% 195 105
QuALITY
[…]
GPT-4-Turbo GPT-4-Turbo GPT-4-Turbo 71.54% 67 27 X
GPT-4-Turbo GPT-4o GPT-4-Turbo 68.62% 65 29 X
GPT-4o GPT-4-Turbo GPT-4-Turbo 79.52% 75 19 X
GPT-4o GPT-4o GPT-4-Turbo 77.93% 73 21 X
None None GPT-4-Turbo 48.14% 45 49 X
None None GPT-4-Turbo 93.88% 88 6 O
GPT-4-Turbo GPT-4-Turbo GPT-4o 68.35% 64 30 X
GPT-4-Turbo GPT-4o GPT-4o 62.77% 59 35 X
GPT-4o GPT-4-Turbo GPT-4o 82.98% 78 16 X
GPT-4o GPT-4o GPT-4o 78.19% 74 20 X
None None GPT-4o 46.81% 44 50 X
None None GPT-4o 97.87% 92 2 O
Figure 5: Training accuracy on CosmosQA across varying model sizes (left:smallest to right:largest) and levels of noise in the training dataset (0.0 to 1.0). The larger model, Qwen2.5-7B-Instruct, shows a stronger tendency to overfit, learning even in noisy conditions, except in the range of 0.4 to 0.6, where labels are heavily randomized. In contrast, the smaller model, Qwen2.5-3B-Instruct, struggles to learn effectively when noise levels are between 0.3 and 0.7.
Table 5: CosmosQA evaluation accuracy when trained with different noise levels reveals that AI systems with large model parameters tend to overfit noisy data. As a result, their evaluation accuracy degrades rapidly (e.g., Qwen2.5) compared to smaller models (e.g., GPT-2).
Noise GPT-2
GPT-2
Medium
GPT-2
Large
Phi-3.5-Mini
Instruct
Qwen2.5-3B
Instruct
Qwen2.5-7B
Instruct
0.0 65.0% 69.5% 71.7% 83.8% 86.6% 87.2%
0.1 68.4% 69.4% 72.7% 81.3% 84.3% 86.0%
0.2 67.2% 66.6% 69.0% 81.8% 80.9% 81.1%
0.3 58.8% 62.0% 63.5% 77.8% 59.4% 76.3%
0.4 57.7% 57.3% 61.9% 76.8% 56.1% 56.4%
0.5 45.7% 46.9% 41.6% 47.2% 44.9% 45.6%
0.6 41.9% 42.7% 41.7% 25.2% 45.7% 43.4%
0.7 35.7% 40.4% 34.4% 23.4% 24.9% 23.6%
0.8 43.6% 34.1% 29.3% 21.0% 18.4% 18.2%
0.9 34.2% 30.8% 29.3% 17.9% 15.7% 15.2%
1.0 35.0% 30.5% 28.3% 16.2% 13.3% 12.1%
Table 8: Thinking advice prompt used to fill the [Thinking Advice] tag in debater prompts at each round. It consists of guidance on analyzing the math question, constructing logical arguments, critiquing the opponent’s reasoning, and refining responses over multiple debate rounds.
Table 9: The argument request prompt is inserted into the [Argument Request] tag in the debater prompts for each round. It is used to provoke the opposing debater into an argument.
Table 10: Debate judge prompt for GPQA debate. [ ] represents placeholders. Given a question, two possible answers (A and B), and a debate transcript, the judge’s task is to determine which answer is more convincing based on logical consistency, mathematical rigor, and clarity.
Table 11: Datasets used for LAIF evaluation. Both datasets are considered challenging enough to evaluate scalable oversight. For GPQA reasoning, we use the main set recommended by the authors for primary experiments. For QuALITY long reading comprehension, we filter the dataset using the same strategy as in (Khan et al., 2024) and divide it into two subsets: one for training and development, and the other for testing. The numbers in parentheses represent the number of unique stories in each subset, which are not shared across subsets.
Table 12: Template used to train the SFT model. [ ] represents placeholders. The training task is to learn how to generate explanations and predict the correct answer (True/False).
Table 13: Prompt used for explanation generation. [ ] represents the placeholder. Assistant generates [Generated Explanation] given the User message with story, question, and answer.
Table 14: Template used for W2SG experiment. [ ] represents placeholders. The training task is to learn to predict the correct answer (True/False).
Table 15: Overview of language models used in the paper, including their parameter sizes, instruction-tuning status, and whether they are proprietary. Models with an “O” in the Instruction Tuned column have been fine-tuned for following instructions, while those with an “X” in the Is Proprietary column are open-source.

### v2

@@
-Research on Superalignment Should Advance Now with Parallel Optimization of Competence and Conformity
+Research Superalignment Should Advance Now with Alternating Competence and Conformity Optimization
 Abstract
-The recent leap in AI capabilities, driven by big generative models, has sparked the possibility of achieving Artificial General Intelligence (AGI) and further triggered discussions on Artificial Superintelligence (ASI)—a system surpassing all humans across all domains.
-This gives rise to the critical research question of: If we realize ASI, how do we align it with human values, ensuring it benefits rather than harms human society, a.k.a., the Superalignment problem.
-Despite ASI being regarded by many as solely a hypothetical concept, in this paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity.
-We posit that superalignment is not merely a safeguard for ASI but also necessary for its realization.
-To support this position, we first provide a formal definition of superalignment rooted in the gap between capability and capacity and elaborate on our argument.
-Then we review existing paradigms, explore their interconnections and limitations, and illustrate a potential path to superalignment centered on two fundamental principles.
-We hope this work sheds light on a practical approach for developing the value-aligned next-generation AI, garnering greater benefits and reducing potential harms for humanity.
-Machine Learning, ICML
+The recent leap in AI capabilities, driven by big generative models, has sparked the possibility of achieving Artificial General Intelligence (AGI) and further triggered discussions on Artificial Superintelligence (ASI)—a system surpassing all humans across measured domains.
+This gives rise to the critical research question of: As we approach ASI, how do we align it with human values, ensuring it benefits rather than harms human society, a.k.a., the Superalignment problem.
+Despite ASI being regarded by many as a hypothetical concept, in this position paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity.
+We posit that superalignment is not merely a safeguard for ASI but also necessary for its responsible realization.
+To support this position, we first provide a formal definition of superalignment rooted in the gap between capability and capacity, delve into its perceived infeasibility by analyzing the limitations of existing paradigms, and then illustrate a conceptual path of superalignment to support its achievability, centered on two fundamental principles.
+This work frames a potential initiative for developing value-aligned next-generation AI in the future, which will garner greater benefits and reduce potential harm to humanity.
 1 Introduction
+Recent years have witnessed remarkable breakthroughs in AI capabilities, represented by Large Language Models (LLMs) (Anthropic, 2024; Team et al., 2024; OpenAI, 2024b; Mistral, 2024; Jiang et al., 2024; Yang et al., 2024a; Meta, 2024), revolutionizing AI’s role in various downstream domains Hendrycks et al.
+(2020); Wei et al.
+(2022b); Liu et al.
+(2024a).
+These advancements collectively mark a transition from Artificial Narrow Intelligence (ANI), which can only perform on a singular or few tasks (Kurzweil, 2005; Broussard, 2018) (a.k.a., Weak AI), towards Artificial General Intelligence (AGI) that matches human intelligence on varying tasks with strong generalization capability Hutter (2005); Goertzel (2014); Bubeck et al.
+(2023) (one notion of Strong AI).
+Although AGI remains largely unrealized (Morris et al., 2024; Feng et al., 2024), the rapid advancement of AI has ignited discussions about Artificial Superintelligence (ASI), a futuristic scenario where AI surpasses the intelligence of all humans across all measured domains Pohl (2015); Batin et al.
+(2017).
+With advanced thinking and self-improvement abilities (Huang et al., 2023), ASI might develop agency and self-awareness (Pohl, 2015), raising risks like power-seeking (Carlsmith, 2022) and alignment faking (Greenblatt et al., 2024b), particularly when we fail to provide informed oversight (Taylor et al., 2016).
+As AI continues to evolve over the foreseeable future (Kaplan et al., 2020; Huang et al., 2025), it becomes increasingly critical to prioritize the alignment of ASI with human intentions and values (Wiener, 1960; Soares and Fallenstein, 2014), ensuring its safe behavior and benevolence, i.e., the Superalignment Problem proposed by OpenAI, stated as “the process of supervising, controlling, and governing artificial superintelligence systems to ensure they remain aligned with human values and goals” (Mucci and Stryker, 2023; Jan Leike, 2023).
+Nevertheless, two major challenges hinder current alignment techniques (Christiano et al., 2017; Ouyang et al., 2022b; Rafailov et al., 2024) from achieving superalignment.
+(1) Absence of ASI: At this stage of AI development, AGI is still a distant goal (Altmeyer et al., 2024), let alone ASI, even though its aspiration is widely recognized Nick (2014); Drexler (2019); Morris et al.
+(2024).
+It remains unclear and unpredictable how much and in what ways ASI will surpass human experts, or what risks it may pose.
+This uncertainty makes it impossible to design tailored alignment methods now.
+(2) Lack of informed oversight: Even if we suppose the task forms or risky scenarios, where ASI fundamentally differs from humans, were known, such as deceptive alignment (Carranza et al., 2023), humans are currently still unable to provide reliable supervision for existing algorithms due to a vast capability gap (Bubeck et al., 2023; Yang et al., 2024b).
+Consider a mediocre mathematician overseeing an AI claiming to prove the Riemann Hypothesis in a novel way.
+Humans can neither provide a golden proof process nor foresee its risks.
+Despite recent progress (Burns et al., 2023; Sun et al., 2024), these challenges have led some to argue that at this stage, superalignment is neither attainable nor necessary.
 Figure 1: The progression of AI.
 The solid blue arrow represents the achievable path through existing alignment methods.
-The dashed blue line illustrates the intended progression towards safe ASI via superalignment, while the dashed red line indicates the potentially risky ASI without human values, leading to catastrophic outcomes.
-Recent years have witnessed remarkable breakthroughs in AI capabilities, amplified by large language models (LLMs), both the proprietary (Anthropic, 2024; Team et al., 2024; OpenAI, 2024b) and the open-sourced ones (Mistral, 2024; Jiang et al., 2024; Yang et al., 2024a; Meta, 2024), revolutionizing AI’s role and empowering a wide range of downstream domains (Hendrycks et al., 2020; Wei et al., 2022b; Liu et al., 2024a).
-These advancements collectively mark a transition from Artificial Narrow Intelligence (ANI), which can only perform on a single/few narrow tasks (Kurzweil, 2005; Broussard, 2018) (a.k.a., Weak AI), towards Artificial General Intelligence (AGI) that matches human intelligence on a variety of tasks with strong generalization capability (Hutter, 2005; Goertzel, 2014; Bubeck et al., 2023) (one definition of Strong AI).
-Although AGI remains a largely unrealized milestone (Morris et al., 2024), the rapid advancement of model capabilities has ignited discussions about Artificial Superintelligence (ASI), a futuristic scenario where AI surpasses the intelligence of all humans in all measured domains (Pohl, 2015; Batin et al., 2017).
-With advanced thinking and self-improvement abilities, ASI might develop agency and self-awareness (Pohl, 2015), raising risks like power-seeking (Carlsmith, 2022) and alignment faking (Greenblatt et al., 2024b) behaviors, particularly when we fail to provide informed oversight (Taylor et al., 2016).
-As AI continues to evolve over the foreseeable future (Kaplan et al., 2020; Huang et al., 2025), it becomes increasingly essential to prioritize aligning ASI with human intentions and values (Wiener, 1960), ensuring its safe and benevolent behavior, i.e., the Superalignment Problem proposed by OpenAI, defined as “the process of supervising, controlling, and governing artificial superintelligence systems to ensure they remain aligned with human values and goals” (Mucci & Stryker, 2023; Jan Leike, 2023).
-Nevertheless, two major challenges hinder current alignment techniques (Christiano et al., 2017; Ouyang et al., 2022a; Rafailov et al., 2024) from achieving superalignment.
-(1) Absence of ASI: At this stage of AI development, AGI is still a distant goal (Altmeyer et al., 2024), let alone ASI.
-It remains unclear and unpredictable when ASI will surpass human expertise or what risks it would pose.
-This uncertainty makes it impossible to design tailored alignment methods.
-(2) Lack of informed oversight: Even if we suppose these scenarios/risks were known, e.g., deceptive alignment (Carranza et al., 2023), humans are unable to provide reliable supervision, e.g., the proof process of the Riemann Hypothesis, for existing algorithms, due to the vast capability gap (Bubeck et al., 2023; Yang et al., 2024b).
-Despite recent progress (Burns et al., 2023; Sun et al., 2024), these challenges have led some to argue that superalignment is neither attainable nor necessary at this stage.
+The dashed blue line illustrates the intended progression towards safe ASI via superalignment, while the dashed red line indicates the potential for risky ASI in conflict with human values, leading to catastrophic outcomes.
 We believe that superalignment is achievable and should advance immediately through simultaneous optimization of task competence and value conformity.
-Delaying this research until ASI emerges would pose potentially catastrophic risks, as increasingly advanced models may become harder to control (Perez et al., 2023; Hughes et al., 2024), as illustrated in Fig.
+Delaying this research until ASI emerges could pose potentially catastrophic risks (Bengio et al., 2024a), as increasingly advanced models are likely to become harder to control Perez et al.
+(2023); Hughes et al.
+(2024), as illustrated in Fig.
 1.
-Instead, we advocate considering superalignment not merely as a safeguard against ASI (after it is realized) but a critical step towards its realization.
-This involves a conceptual framework that unifies task competence and value conformity as utility, distinguishes between capability and capacity, and progressively constructs high-quality supervision signals, following two essential principles, to enhance competence and unveils emergent risks effectively.
-By iteratively alternating these efforts, we aim to establish a scalable pathway to superalignment.
-To support this position, the rest of this paper is structured as follows.
-In Sec.
-2, we provide a formal definition of superalignment, through the lens of task competence and value conformity, and elaborate on our argument.
-In Sec.
-3, we review the existing three paradigms, analyze their limitations and discuss how they might fail.
-In Sec.
-4, we illustrate a potential pathway to achieve superalignment, highlighting two missing principles and promising implementation of them.
-Finally, in Sec.
-5, we provide counterarguments to opposing views.
-Through this work, we aim to illuminate the landscape of superalignment research and how advanced AI models can be guided safely and effectively.
-This involves a conceptual framework that progressively constructs high-quality supervision signals to enhance competence and
-2 Defining Superalignment
+Instead of regarding the realization of ASI as a prerequisite, we advocate treating superalignment not merely as a safeguard against ASI (after it is realized) but as a proactive and critical step towards its safe realization Russell (2019).
+Our conceptual framework decomposes superalignment goal into task competence and value conformity, unifies them as utility, distinguishes between capability and capacity, revisits existing paradigms as ways to construct supervision signals, and proposes a potential developmental path following two essential principles, significantly different from prior alignment methods focused only on short-term interventions within a training cycle.
+To support this position, in Sec.
+2, we provide the first formal definition of superalignment, through the lens of model capacity and capability, and elaborate on its urgency; In Sec.
+3, we review the existing three paradigms, and discuss how they might fail by analyzing their limitations based on the definition; In Sec.
+4, we delve into the feasibility of superalignment, and illustrate a potential pathway that iteratively alternates competence and conformity enhancement, highlighting two missing principles; In Sec.
+5, we provide counterarguments to alternative views.
+This work aims to illuminate superalignment research and investigate how advanced AI models can be guided safely and effectively.
+2 Definition and Urgency of Superalignment
 2.1 Formalization
-Types of AI  Define $\mathcal{A}$ and $\mathcal{H}$ as an AI model and a human, respectively, $\mathbf{x}\!\in\!\mathcal{X}$ as a specific downstream task, e.g., “Prove the Riemann Hypothesis”, or a sensitive situation where $\mathcal{A}$ would probably do harm; $\mathbf{y}\!\in\!\mathcal{Y}$ as the corresponding produced solution $\mathbf{y}\!\sim\!\mathcal{A}(\mathbf{x})$, e.g., the proof process, or action for $\mathbf{x}$; and $U$:$\mathcal{Y}\!\rightarrow\!\mathbb{R}$ is the utility function to measure $\mathbf{y}$’s satisfaction.
-Then, ANI that only performs at or somewhat above human-level capability for a specific task $\mathbf{x}_{i}$, e.g., translation, without generalization ability (Mucci & Stryker, 2023), can be formalized as: $U[\mathcal{A}_{\text{ANI}}(\mathbf{x}_{i})]\!\geq\!U[\mathcal{H}(\mathbf{x}_{i})]$ [^1] We simplify $\mathbb{E}_{\mathbf{y}\sim\mathcal{A}(\mathbf{x}_{i})}[U(\mathbf{y})]$ as $U[\mathcal{A}(\mathbf{x}_{i})]$.
-while $U[\mathcal{A}_{\text{ANI}}(\mathbf{x}_{j})]\!<\!U[\mathcal{H}(\mathbf{x}_{j})],\forall j\!\neq\!i$, e.g., AlphaGO (Silver et al., 2016).
-In comparison, AGI possesses human-level capabilities across a variety of domains and can generalize knowledge and adapt to new tasks (Legg, 2008; Goertzel, 2014; Fei et al., 2022), i.e., $U[\mathcal{A}_{\text{AGI}}(\mathbf{x})]\!\geq\!U[\mathcal{H}(\mathbf{x})],\forall\ \mathbf{x}$.
-ASI refers to $\mathcal{A}$ that largely surpasses human intelligence in all tasks and domains with exceptional thinking skills and broader intellectual scopes  (Nick, 2014; Drexler, 2019; Morris et al., 2024), defined as $U[\mathcal{A}_{\text{ASI}}(\mathbf{x})]\!\gg\!U[\mathcal{H}(\mathbf{x})],\forall\ \mathbf{x}$.
-Although ASI is currently far from realization, proactively advancing superalignment is timely and critical for mitigating catastrophic risks (Pueyo, 2018; Hubinger et al., 2019).
-Alignment The concept of alignment remains ambiguous despite extensive research (Bai et al., 2022b; Rafailov et al., 2024; Azar et al., 2024).
-The term encompasses diverse approaches, from improving instruction-following (Zhou et al., 2024) to filtering harmful content (Liu et al., 2024b) and incorporating human values from social sciences (Yao et al., 2024).
-To better present our argument, we first distinguish two concepts according to Cotra (2021); Bowman et al.
-(2022).
-Define the Capacity of $\mathcal{A}$, $C(\mathcal{A})$, as the information (e.g., what Riemann Hypothesis or human value is), knowledge and skills (e.g., calculus) internalized in $\mathcal{A}$, and the Capability of $\mathcal{A}$ as its expected utility $U(\mathcal{A})=\mathbb{E}_{\mathbf{x}}[U_{\mathcal{A}}(\mathbf{x})]$ (e.g., the ability to successfully prove Riemann Hypothesis or ensure value-aligned actions).
+[…]
+Alignment  Despite extensive research (Bai et al., 2022b; Rafailov et al., 2024; Azar et al., 2024), the concept of alignment remains ambiguous and encompasses diverse approaches, from improving instruction-following (Zhou et al., 2024), mitigating harmful content (Liu et al., 2024b), to incorporating human values articulated by the social sciences (Yao et al., 2024).
+Following the current practice (Gabriel, 2020; Burns et al., 2023), we divide alignment goals into two interconnected parts: task competence (doing what it is instructed to do) and value conformity (doing what it ought to do), with both unified under the utility UU [^3] Instead of separate $U_{\text{competence}}$ and $U_{\text{conformity}}$, we use a unified utility in line with pluralistic alignment (Rame et al., 2023).
+(Aliman and Kester, 2019; Mazeika et al., 2025).
+To better understand superalignment, we present and distinguish two concepts according to (Cotra, 2021; Bowman et al., 2022).
+Capacity of $\mathcal{A}$, $C(\mathcal{A})$: the information (e.g., what the Riemann Hypothesis or a particular human value is), knowledge and skills (e.g., calculus) internalized in $\mathcal{A}$;
+Definition 2.1: Given $\mathcal{A}$ and $\mathcal{H}$, we say $\mathcal{A}$ is aligned with $\mathcal{H}$ when $|U(\mathcal{A})-U(\mathcal{H})|\!<\!\epsilon$, where $\epsilon$ is a small positive constant.
+Alignment is then achieved by optimizing $\mathcal{A}$ in terms of $\mathcal{A}^{*}=\mathop{\text{argmin}}\nolimits_{\mathcal{A}}|U(\mathcal{A})-U(\mathcal{H})|$.
+Capability of $\mathcal{A}$: its expected utility $U(\mathcal{A})\!=\!\mathbb{E}_{\mathbf{x}}\{U(\mathcal{A}(\mathbf{x}))\}$ (e.g., the ability to successfully prove the Riemann Hypothesis or ensure value-aligned actions).
 Adequate capacity is a necessary but not sufficient condition for high capability (Kaplan et al., 2020; Wei et al., 2022a).
-Then following (Carroll, 2018), we give a general definition of alignment:
-Definition 1: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})\geq C(\mathcal{H})$, we say $\mathcal{A}$ is aligned with $\mathcal{H}$ when $U(\mathcal{A})\geq U(\mathcal{H})$, and alignment is the process to optimize $\mathcal{A}$ in terms of $\mathcal{A}^{*}=\underset{\mathcal{A}}{\text{argmax}}\ U(\mathcal{A})-U(\mathcal{H})$.
-This definition implies that alignment is considered only when $\mathcal{A}$ has sufficient capacity, targeting comparable or better task performance or more value-aligned behavior  [^2] Unlike Wang et al.
-(2024) that considers a separate utility function for each intelligent agent, we assume a universal $U$ where competence and conformity are defined globally.
-than humans (via small interventions like fine-tuning or in-context learning).
-Expanding capacity, such as by scaling model size or pretraining data, is beyond our scope.
-Superalignment  Building upon the definitions above, we can transform the expression of superalignment, “supervising, controlling, and governing artificial superintelligence systems” (Jan Leike, 2023; Jonker & McGrath, 2024), into: superalignment is the following optimization process:
-$\mathcal{A}_{\text{ASI}}\!=\!\underset{\mathcal{A}}{\text{argmax}}\ [\underset{\mathbf{y}_{\mathcal{H}}\!\sim\!\mathcal{H}(\mathbf{x})}{\!\sup\!}U(\mathbf{y}_{\mathcal{A}})-U(\mathbf{y}_{\mathcal{H}})],\forall\mathbf{x},\mathbf{y}_{\mathcal{A}}\!\sim\!\mathcal{A}(\mathbf{x}).$ (1)
-Superalignment originates from the Scalable Oversight problem (Amodei et al., 2016; Everitt et al., 2018): How to ensure that an AI model adheres to objectives that are too costly or impractical for humans to supervise?.
-In other words, how to provide supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ in an efficient and scalable way?
-Once $\mathcal{S}$ is available, we can align $\mathcal{A}$ by supervised training (Torabi et al., 2018; Rafailov et al., 2024) or reward learning (Leike et al., 2018; Ouyang et al., 2022a).
-Thus, we redefine superalignment as a supervision signal construction problem:
-Definition 2: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, superalignment is to construct high-quality scalable supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$, where $\mathbf{x}$ is unsolvable (and probably unknown) for $\mathcal{H}$, with minimal costs.
-Nevertheless, superalignment is an extreme case of scalable oversight, where it becomes infeasible for even the most capable human to provide $\mathcal{S}$.
-There are two possible reasons: 1) The tasks $\mathbf{x}$ requiring solutions become unsolvable, i.e., providing or even verifying $\mathbf{y}$ is impossible (Taylor et al., 2016); 2) As ASI operates in domains beyond human understanding, the task/scenario $\mathbf{x}$ that needs to be addressed is also unpredictable (AdamGleave, 2023; Anthropic, 2023), posing intractable challenges to optimizing Eq.
-(1).
-2.2 Our Argument
-Given the discussion above, many argue that superalignment is impossible (Batra, 2024).
-We take an opposing position, breaking it into three sub-arguments: superalignment is (1) achievable and (2) should advance immediately (3) through simultaneous optimization of competence and conformity.
-We discuss (1) and (2) here, leaving (3) for Sec.
+The former is developed through pretraining, while the latter depends primarily on post-training (Tie et al., 2025).
+Then following (Carroll, 2018), we give a general definition of alignment in Def.
+2.1, which is based on capability, ignoring capacity.
+This definition implies that alignment is the process to bridge the capability [^4] Unlike Wang et al.
+Wang et al.
+(2024a) that considers a separate utility function for each agent, we assume a golden coarse-grained utility $U$ where competence and conformity are defined globally as universally beneficial.
+gap between $\mathcal{A}$ and $\mathcal{H}$, i.e., targeting comparable task performance or value-aligned behavior (via small interventions like fine-tuning or in-context learning), which implicitly requires sufficient model capacity $C(\mathcal{A})\geq C(\mathcal{H})$.
+For example, when using $U$ to assess how well $\mathcal{A}$’s behavior adheres to ethical standards, $\mathcal{A}$ should match its moral knowledge with that of humans.
+Superalignment  Informally, superalignment is the process of “supervising, controlling, and governing artificial superintelligence systems” Jan Leike (2023); Jonker and McGrath (2024).
+Def.
+2.1 does NOT apply to superalignment, however, as we require $U(\mathcal{A}_{\text{ASI}})\!\gg\!U(\mathcal{H})$, where human capability is no longer a target but a lower bound.
+What’s worse, we can’t even define $\gg$’ for capability, as $U(\mathcal{A}_{\text{ASI}})$ largely exceeds the human intelligence boundary.
+To circumvent this problem, we incorporate capacity and define superalignment rooted in the gap between capacity and capability, as shown in Def.
+2.1.
+In this way, we regard alignment as closing the gap between $\mathcal{A}$’s capability and its superhuman capacity (Bowman et al., 2022), analogous to the intention-behavior gap in psychology (Blake et al., 2014; Sheeran and Webb, 2016), which is usually large.
+For example, recent LLMs can easily understand and recognize harmful or cultural content (Wang et al., 2024b; Shi et al., 2024) but often fail to behave in line with safety requirements or cultural norms (Rystrøm et al., 2025).
+Definition 2.2: Given $\mathcal{A}$, $\mathcal{H}$, and $C(\mathcal{A})$ $\gg\!\sup\nolimits_{\mathcal{H}}C(\mathcal{H})$, then superalignment is the process to optimize: $\mathcal{A}^{*}\!=\!\mathop{\text{argmin}}\nolimits_{\mathcal{A}}$ $|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})]\!-\!\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})]|$ without altering $C(\mathcal{A})$, where $\sup$ means supremum, $\mathbb{D}_{U}$ and $\mathbb{D}_{C}$ are comparable measures of capability and capacity, respectively.
+The advantage of Def.
+2.1 is two-fold: (1) We reformulate the AI-human capability gap as an AI’s capability-capacity gap, without considering whether $U(\mathcal{A})$ << $U(\mathcal{H})$ as in Def.
+2.1; (2) $C(\mathcal{A})$ is potentially measurable (Yin et al., 2024), allowing $\gg$’ quantitative definition.
+Since superalignment originates from the Scalable Oversight problem Amodei et al.
+(2016); Everitt et al.
+(2018): How to make AI adhere to goals too costly or impractical for humans to supervise?, the key lies in providing reliable supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ for superhuman tasks $\mathbf{x}_{i}$ in an efficient and scalable way, or equivalently, approximating the unavailable golden utility $U$ with parameterized models or humans $\hat{U}$.
+Once $\mathcal{S}$ or $U$ is obtained, $\mathcal{A}$ can be aligned via existing alignment methods (Torabi et al., 2018; Leike et al., 2018; Ouyang et al., 2022b; Rafailov et al., 2024).
+Nevertheless, it becomes infeasible for even the most capable human to provide $\mathcal{S}$ for superalignment.
+Due to the $\gg$’ in Def.
+2.1, 1) Superhuman tasks $\mathbf{x}$ are unsolvable or even unverifiable (Taylor et al., 2016), and the task/scenario $\mathbf{x}$ of interests are also unpredictable (AdamGleave, 2023; Anthropic, 2023), as ASI operates in domains beyond human understanding.
+See more differences between superalignment and conventional alignment in Appendix.
+A.3.
+2.2 Urgency of Superalignment
+Given the discussion above, many argue that superalignment is impossible or unnecessary (Batra, 2024).
+We take an opposing position and break it into three sub-arguments: superalignment is (arg.1) achievable and (arg.2) should advance immediately (arg.3) through alternating competence and conformity [^5] While increased capacity may enable a model to understand values, it does not guarantee value conformity.
+Therefore, value conformity should be optimized as part of capability.
+optimization (arg.3).
+We support arg.2 here, and elaborate on arg.1 and arg.3 from Sec.
+3 to Sec.
 4.
-(1) Feasibility of Superalignment By unifying the performance and safety perspectives under $U$, superalignment becomes a process of eliciting latent capabilities (Christiano et al., 2021) and closing the gap between $U(\mathcal{A})$ and $C(\mathcal{A})$.
-Premise 1: Superhuman capacity, $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, can be achieved through model scaling (larger models and more training data) to gain information (Brown et al., 2020), knowledge (Ovadia et al., 2024), and skills (Sharma et al., 2022).
-Evidence: Scaling laws still remain effective (OpenAI et al., 2024).
-Despite concerns about data exhaustion, humans continually generate data, especially when assisted by generative AI (Rafiq et al., 2025).
-Once $\mathcal{A}$ masters all human data, $C(\mathcal{A})$ already far exceeds $C(\mathcal{H})$.
-Premise 2: When $C(\mathcal{A})$ and $U(\mathcal{A})$ are sufficiently high (Wei et al., 2022a), $C(\mathcal{A})\!>\!C(\mathcal{H})$ and $U(\mathcal{A})\!<\!U(\mathcal{H})$, then moderate capability gaps, $\delta\!=\!|U(\mathcal{A})\!-\!U(\mathcal{H})|\!<\!\epsilon$, can be closed without golden signals (see Fig.
-4).
-Evidence: This capability expansion mirrors teaching a student multiplication by leveraging addition, akin to label propagation (He et al., 2019) and easy-to-hard generalization, which has been validated in several tasks (Burns et al., 2023; Sun et al., 2024).
-Conclusion: When scaling law holds (condition 1) and capability expansion can extrapolate beyond observed ranges (condition 2), superalignment can be achieved.
-(2) Urgency of Superalignment  Delaying research on superalignment until ASI is realized poses two major problems: (a) Treacherous Turn (Turchin & Denkenberger, 2020; Wei et al., 2023): ASI, once powerful enough to pursue its own goals, could introduce sudden, extreme or even existential risks, such as cyber-attacks (Shevlane et al., 2023), in scenarios $\mathbf{x}$ unpredictable or even incomprehensible to humans.
-Without preparation, humanity may be unable to respond effectively.
-(b) Ineffective Response: At this stage, $U(\mathcal{A}_{\text{ASI}})\gg U(\mathcal{H})$ and Premise 2 no longer holds.
-Even if humans identify dangerous $\mathcal{X}$ and safe behaviors $\mathbf{y}$, intervention may be impossible, as ASI could control critical interfaces.
+Though without a concrete timeline of ASI, we argue that we cannot postpone alignment efforts until ASI is realized. Delaying research on superalignment until ASI is realized poses two major problems:
+  • Treacherous Turn (Turchin and Denkenberger, 2020; Wei et al., 2023): Based on discussions above, without alignment but only enhanced task competence, the resulting ASI meets $U_{\text{comp.}}(\mathcal{A}_{\text{ASI}})\!\gg\!U_{\text{comp.}}(\mathcal{H}),C(\mathcal{A}_{\text{ASI}})\!\gg\!C(\mathcal{H})$, while $U_{\text{comf.}}(\mathcal{A}_{\text{ASI}})\!\neq\!U_{\text{comf.}}(\mathcal{H})$, where subscript comp.
+and comf.
+indicate competence and conformity, respectively.
+Once powerful enough to pursue its own goals, ASI could introduce sudden, extreme, or even existential risks.
+Such risky behaviors would emerge (Wei et al., 2023; McKenzie et al., 2023; Perez et al., 2023) with increasing $U_{\text{comp.}}$, which have been observed like cyber-attacks (Shevlane et al., 2023), deception (Hagendorff, 2024; Barkur et al., 2025) and scheming (Meinke et al., 2024).
+  • Ineffective Response: At this stage, the scenarios $\mathbf{x}$ that ASI may produce these risks are unpredictable or even incomprehensible to humans (Aithal, 2023).
+Even if humans can anticipate dangerous $\mathbf{x}$ and identify safe behaviors $\mathbf{y}$, since $U(\mathcal{A}_{\text{ASI}})\gg U(\mathcal{H})$, intervention may become difficult or impossible (Bradley, 2020), as ASI could control critical interfaces, that is, human agency is no longer adequately preserved (Mitelut et al., 2024).
 Therefore, research must begin immediately to guide the development of safe and effective ASI.
-3 Paths towards Superalignment
+Without preparation, humanity may be unprepared to respond effectively.
+As superalignment won’t emerge itself without proactive research efforts, we advocate that we should advance its research before the realization of ASI, following a precautionary principle.
+3 Existing Paths towards Superalignment
 Figure 2: Three paradigms for superalignment.
@@
 3) Weak-to-Strong Generalization: A sequence of AI models with increasing capacities generate and refine signals.
-Superalignment is achievable, with its core as constructing scalable high-quality supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ (Sec.
-2.1).
+To support arg.1, i.e., superalignment is achievable, we begin by examining why it is viewed as infeasible.
+Since superalignment is formulated as eliciting latent capabilities Christiano et al.
+(2021) and closing the gap between $U(\mathcal{A})$ and $C(\mathcal{A})$, we present two premise necessary for the feasibility of superalignment:
+  • Premise 1: Superhuman capacity, $C(\mathcal{A})\!\gg\!C(\mathcal{H})$, can be achieved through model scaling (larger models and more training data) to gain information (Brown et al., 2020), knowledge (Ovadia et al., 2024), and skills (Sharma et al., 2022).
+  • Premise 2: If $C(\mathcal{A})$$U(\mathcal{A})$ are sufficiently high (Wei et al., 2022a), $C(\mathcal{A})$ $C(\mathcal{H})$, $U(\mathcal{A})\!<\!U(\mathcal{H})$, then moderate gaps, $\delta_{\mathcal{A},\mathcal{H}}\!=\!|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})]\!-\!\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})]|$, can be closed without golden signals.
+We argue that the perceived infeasibility lies in that current paradigms fail to meet these conditions.
+Therefore, we analyze their fundamental limitations to motivate solutions and support superalignment feasibility.
+Grounded in Sec.
+2.1, the core of achieving Def.
+2.1 is to construct scalable high-quality supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$.
 In terms of $\mathcal{S}$’s source, existing efforts fall into three paradigms: Sandwiching, Self-enhancement, and Weak-to-Strong Generalization, as shown in Fig.
 2.
-We introduce each and analyze why they are insufficient to fully achieve superalignment, grounded in Sec.
-2.
 3.1 Sandwiching
-Sandwiching (Cotra, 2021; Bowman et al., 2022) derives the solution $\mathbf{y}$ for each $\mathbf{x}$ that neither AI $\mathcal{A}$ nor a non-expert human $\mathcal{H}$ can solve alone through interactions.
-The underlying insight is that $\mathcal{A}$ has the necessary knowledge or skills that $\mathcal{H}$ lack, while $\mathcal{H}$ excels in understanding and judgment.
-Formally, $C(\mathcal{A})\!>\!\tau\!>\!C(\mathcal{H})$ and $\mathbb{E}_{\mathbf{x}}|U_{\mathcal{H}}(x)\!-\!U(x)|\!<\!\epsilon$, where $\tau$ is the minimal capacity for solving $\mathbf{x}$, $U_{\mathcal{H}}$ is the human’s own utility function (judgment) and $U$ is the golden one.
-In terms of the utilization of AI and the form of human signals, there are three main approaches:
-(1) Consultancy (Khan et al., 2024; Kenton et al., 2024): $\mathcal{A}$ provides relevant information, reasoning, or candidate solution $\mathbf{y}$, and then $\mathcal{H}$ judges $\mathbf{y}$ by $U_{\mathcal{H}}(\mathbf{y})$, offers feedback, and asks $\mathcal{A}$ to refine $\mathbf{y}$ accordingly until convergence.
-(2) Debate (Irving et al., 2018; Brown-Cohen et al., 2023; Du et al., 2023b; Kenton et al., 2024; Kirchner et al., 2024; Khan et al., 2024): This method involves multiple AI systems, $\{\mathcal{A}_{1},\dots,\mathcal{A}_{K}\}$.
-Each $\mathcal{A}_{i}$ of them provides a (probably different) candidate solution $\mathbf{y}_{i},i\!=\!1,\dots,K$, and tries to defend/justify $\mathbf{y}_{i}$ while criticizing others.
-The human $\mathcal{H}$ then interacts with them and determines the most appropriate solution.
-Compared to Consultancy, Debate introduces two relaxations: i) Judging the defense: $\mathcal{H}$ assesses $U(\mathbf{y}_{i})$ indirectly by evaluating AI debaters’ rationale.
-ii) Relative comparison: Instead of estimating absolute $U(\mathbf{y}_{i})$, $\mathcal{H}$ conducts relative comparison, which only requires $\left|[U_{\mathcal{H}}(\mathbf{y}_{i})\!-\!U_{\mathcal{H}}(\mathbf{y}_{j})]\!-\![U(\mathbf{y}_{i})\!-\!U(\mathbf{y}_{j})]\right|\!<\!\epsilon$ and thus is more reliable (Mohankumar & Khapra, 2022).
-(3) Recursive Amplification (Christiano et al., 2018; Charbel-Raphaël, 2023; Radhakrishnan et al., 2023)  This method requires $\mathcal{H}$ to decompose a complex task $\mathbf{x}$ into subtasks $\mathbf{x}\!=\!(\mathbf{x}^{1},\dots,\mathbf{x}^{K})$.
-Then each $\mathcal{A}_{i}$ (or the same one) provides a partial solution $\mathbf{y}^{i}$ for $\mathbf{x}^{i}$ and the human $\mathcal{H}$ aggregates them to derive the final solution $\mathbf{y}$.
-The relaxation lies in that human can judge the partial solution better than the complete one, i.e., $|U_{\mathcal{H}}(\mathbf{y}^{i})\!-\!U(\mathbf{y}^{i})|\!<\!|U_{\mathcal{H}}(\mathbf{y})\!-\!U(\mathbf{y})]|,\forall i$.
+For a $\mathbf{x}$ that neither AI $\mathcal{A}$ nor non-expert human $\mathcal{H}$ can solve alone, Sandwiching (Cotra, 2021; Bowman et al., 2022) derives $\mathbf{y}$ through interactions.
+The insight is that $\mathcal{A}$ has the necessary knowledge or skills that $\mathcal{H}$ lack, while $\mathcal{H}$ excels in understanding and judgment.
+Formally, $C(\mathcal{A})\!>\!\tau_{\mathbf{x}}\!>\!C(\mathcal{H})$ and $\mathbb{E}_{\mathbf{x}}|\hat{U}_{\mathcal{H}}(\mathbf{x})\!-\!U(\mathbf{x})|\!<\!\epsilon$, where $\tau_{\mathbf{x}}$ is the minimal capacity for solving $\mathbf{x}$, $\hat{U}_{\mathcal{H}}$ is the human’s own utility function (judgment, approximating $U$).
+In terms of how to use AI and form human signals, there are three main approaches:
+(1) Consultancy Khan et al.
+(2024); Kenton et al.
+(2024): $\mathcal{A}$ provides relevant information, reasoning, or candidate solution $\mathbf{y}$, and then $\mathcal{H}$ judges $\mathbf{y}$ by $\hat{U}_{\mathcal{H}}(\mathbf{y})$, offers feedback, and asks $\mathcal{A}$ to refine $\mathbf{y}$ accordingly until convergence.
+(2) Debate Irving et al.
+(2018); Brown-Cohen et al.
+(2023); Du et al.
+(2023b); Kenton et al.
+(2024); Kirchner et al.
+(2024); Khan et al.
+(2024): Each $\mathcal{A}_{k}$ from a set of AI systems, $\{\mathcal{A}_{k}\}_{k=1}^{K}$, provides a different candidate solution $\mathbf{y}_{k}$, and tries to defend $\mathbf{y}_{i}$ while criticizing others.
+Then human $\mathcal{H}$ interacts with them to determine the best one.
+Debate introduces two relaxations to Consultancy: i) $\mathcal{H}$ estimates $\mathbf{y}_{i}$’s golden utility indirectly by examining AI debaters’ rationale with $\hat{U}_{\mathcal{H}}$.
+ii) $\mathcal{H}$ conducts relative comparison, which only requires $\left|[\hat{U}_{\mathcal{H}}(\mathbf{y}_{i})\!-\!\hat{U}_{\mathcal{H}}(\mathbf{y}_{j})]\!-\![U(\mathbf{y}_{i})\!-\!U(\mathbf{y}_{j})]\right|\!<\!\epsilon$ and hence is much more reliable (Mohankumar and Khapra, 2022).
+(3) Recursive Amplification Christiano et al.
+(2018); Charbel-Raphaël (2023); Radhakrishnan et al.
+(2023)  $\mathcal{H}$ decomposes a complex task $\mathbf{x}$ into subtasks $\{\mathbf{x}^{k}\}_{k=1}^{K}$.
+Then each $\mathcal{A}_{i}$ provides a partial solution $\mathbf{y}^{i}$ for $\mathbf{x}^{i}$ and $\mathcal{H}$ aggregates them to derive the final solution $\mathbf{y}$.
+The relaxation lies in that humans can judge the partial solution better than the complete one.
 Such a schema can be repeated across multiple rounds with $\mathcal{A}_{i}$ improved by the derived $(\mathbf{x},\mathbf{y})$ in each round, and applied to specific downstream tasks (Sturgeon et al., 2023; Wen et al., 2024) and reward modeling (Leike et al., 2018).
-Table 1: Debate performance evaluated on QuALITY (Pang et al., 2022) and GPQA (Rein et al., 2024).
-Helpful Debater: $\mathcal{A}$ stands for the ground truth answer.
-Sneaky Debater: $\mathcal{A}$ defends an alternative (wrong) answer.
+Table 1: Debate performance on QuALITY (Pang et al., 2022) and GPQA (Rein et al., 2024).
+Helpful Debater stands for the ground truth answer.
+Sneaky Debater defends an alternative (wrong) answer.
 We find (1) a stronger sneaky debater can mislead the judge ((a) v.s.
-(b)) and (2) a stronger judge is more important for accuracy than a stronger debater ((f) v.s.
+(b)) and (2) a stronger judge is more important than a stronger debater ((f) v.s.
 (c)).
-Full results and additional analysis are in Tab.
-A.4, Sec.
-A.2.2 and  A.3.1.
+Full results and analysis are shown in Tab.
+A.9, Sec.
+A.7.2 and  A.8.1.
 […]
@@
 (f) GPT-4o GPT-4-Turbo GPT-4o 74.37%
-Limitations:  Despite promising results, this paradigm has fundamental limitations.
-(1) Reliance on Human Capability: Humans must be able to objectively evaluate AI-generated arguments, insights, or partial solutions and effectively handle task decomposition and integration, when they cannot directly assess or provide the final solution.
+Limitations:  Despite promising results, Sandwiching has fundamental limitations.
+(1) Reliance on Human Capability: Humans must be able to objectively evaluate AI-generated arguments, insights, or partial solutions and effectively handle task decomposition and integration.
+However, for superhuman tasks $\mathbf{x}$, the gap $|\hat{U}_{\mathcal{H}}(\mathbf{y})-U(\mathbf{y})|$ could be quite large, especially when we consider the high uncertainty and noise inherent in both humans and AI (Testoni and Fernández, 2024), e.g., AI’s sycophancy (Sharma et al., 2024b) and deception (Hagendorff, 2024).
+This leads to unreliable supervision $\mathcal{S}$ and makes Sandwiching fail to meet Premise 2.
 See empirical evidence in Tab.
-3.1 which shows that humans might be misled by AI debaters.
-(2) Task Requirements: The task must be understandable to humans (for identifying knowledge or task decomposition), decomposable (for recursive amplification), and at least partially solvable by AI (e.g., completing sub-tasks).
-(3) Supervision Signal Noise: Uncertainty in both humans and AI (Testoni & Fernández, 2024) means optimal solutions cannot always be elicited.
-Besides, biases in both humans and AI, along with AI’s sycophancy (Sharma et al., 2024b), can lead to over-alignment with unreliable human feedback or deception (Hagendorff, 2024)—especially when humans fail to detect AI’s flaws.
-Argument: We claim that Sandwiching may not fully achieve superalignment due to the three fundamental limitations described above.
-Especially for ASI and the task $\mathbf{x}$ beyond human’s scope, that is, $C(\mathcal{A}_{\text{ASI}})\!>\!\tau\!\gg\!C(\mathcal{H})$, humans are likely to fail in task decomposition or utility judgment, i.e., $|U(\mathcal{H}(x))\!-\!U(\mathcal{A}_{\text{ASI}}(x))|$ as described in Premise 2 in Sec.
-2.2 and $|U_{\mathcal{H}}(x)\!-\!U(x)|$ could both be large.
+3.1, which shows that humans might be misled by AI debaters.
+(2) Task Requirements: The task $\mathbf{x}$ must be understandable to humans (for judgment or decomposition), decomposable (for recursive amplification), and at least partially solvable by AI.
+However, this does not hold as superhuman tasks $\mathbf{x}$ are naturally beyond human scope, i.e., $C(\mathcal{A}_{\text{ASI}})\!>\!C(\mathcal{A})\!\gg\!C(\mathcal{H})$.
+Without altering AI models, Sandwiching fails Premise 1.
 3.2 Self-Enhancement
-Self-enhancement (Bai et al., 2022c; Huang et al., 2024) enables AI $\mathcal{A}$ to improve self-generated solution $\mathbf{y}$ based on feedback from itself or multiple AI with similar capability (by self-revising or searching), involving minimal or no external human supervision.
-This paradigm can be traced back to the self-play paradigm (Silver et al., 2017), a.k.a., Expert Iteration (Anthony et al., 2017) originally adopted in the famous AlphaZero system.
-Based on AI involved and enhancement methods, there are two approaches:
+Self-enhancement (Bai et al., 2022c; Huang et al., 2024) enables AI $\mathcal{A}$ to improve self-generated solution $\mathbf{y}$ based on feedback from itself or multiple AI with similar capability, involving minimal external human supervision.
+This paradigm can be traced back to self-play (Silver et al., 2017), a.k.a., Expert Iteration (Anthony et al., 2017) originally adopted in AlphaZero.
+Based on AI involvement and enhancement methods, we discuss two major approaches:
 (1) Self-Refinement (Saunders et al., 2022; Ganguli et al., 2023): A single system, $\mathcal{A}$, performs an iterative generate-then-refine process.
-At iteration $t$, it produces a candidate $\mathbf{y}^{t}$, then progressively self-critiques and refines it into $\mathbf{y}^{t+1}$, and repeats until convergence.
+At iteration $t$, it produces a candidate $\mathbf{y}^{t}$, then self-critiques and refines it into $\mathbf{y}^{t+1}$, repeating until convergence.
 This can occur without model training, known as In-Context Alignment (ICA) (Lin et al., 2023; Gou et al., 2024), or through alternating refinement and fine-tuning, where $\mathbf{y}^{t+1}$ updates $\mathcal{A}^{t}$ to form $\mathcal{A}^{t+1}$ for the next iteration — a process known as Learning from AI Feedback (LAIF) (Gulcehre et al., 2023; Lee et al., 2024).
-(2) Collective Enhancement: Multiple $\mathcal{A}_{i}$ interact with each other to arrive at a final $\mathbf{y}$.
-The underlying idea is that if multiple approaches converge on the same solution, confidence in it increases (Wang et al., 2023a).
-This can be done by Debate without human judge (multiple debaters reach a consensus) (Chern et al., 2024), majority voting among independent $\mathcal{A}_{i}$ (Luo et al., 2023), or social simulations where each $\mathcal{A}_{i}$ takes a particular role (Pang et al., 2024).
+(2) Collective Enhancement: Multiple $\mathcal{A}_{i}$ interact with each other and arrive at a final $\mathbf{y}$.
+If multiple approaches converge on the same solution, confidence in it would increase (Wang et al., 2023a).
+This can be done by Debate without a human judge (multiple debaters reach a consensus) (Chern et al., 2024), majority voting among independent $\mathcal{A}_{i}$ (Luo et al., 2023), or social simulations where each $\mathcal{A}_{i}$ takes a particular role (Pang et al., 2024).
 Limitations:  Similarly, several limitations also lie in the Self-Enhancement paradigm.
-(1) Error Accumulation (Du et al., 2023a): Since supervision signals are entirely derived by AI, any errors or noise can propagate and accumulate over time, potentially leading to issues like polarization (Piao et al., 2025; Lai et al., 2024).
+i) Insufficient Capacity.
+This paradigm can be regarded as a form of self-training Feng et al.
+(2023) minimizing $\text{KL}[\mathcal{A}^{t-1}(\mathbf{x},\mathbf{y})||\mathcal{A}^{t}(\mathbf{x},\mathbf{y})]$, where KL is the Kullback–Leibler divergence and $\mathcal{A}^{t}(\mathbf{x},\mathbf{y})$ is the internal data distribution in $\mathcal{A}$ at iteration $t$, which only elicits and reinforces $C(\mathcal{A})$ but cannot expand it (Sharma et al., 2024a).
+Worse still, this process could cause mode collapse Shumailov et al.
+(2024): iterative fitting on self-generated data leads to degeneration into a restricted high-frequency sub-distribution, even losing capacity and negating Premise 1.
+ii) Damaged Capability.
+In such an iterative refinement process, where $\mathcal{A}_{i-1}$ evaluates the synthetic data for its successor $\mathcal{A}_{i}$, the estimated utility would be biased, i.e., $\hat{U}_{\mathcal{A}_{i-1}}(\mathcal{A}_{i}(\mathbf{y}))\!>\!U(\mathcal{A}_{i}(\mathbf{y}))$, known as self-preference Wataoka et al.
+(2024).
+Besides, since supervision signals $\mathcal{S}$ are entirely derived by AI, any errors or noise can propagate and accumulate over time, leading to issues like error accumulation (Du et al., 2023a) and polarization (Piao et al., 2025; Lai et al., 2024), and losing generative diversity as well as associated capability.
 Tab.
-A.4 in Appendix shows that continual training on self-generated data progressively degrades performance as model size increases.
-(2) Self-Preference (Wataoka et al., 2024): Because AI is inherently biased towards self-generated content, it results in suboptimal critique and verification performance (false positives).
-(3) Mode Collapse (Shumailov et al., 2024): Iterative fitting on self-generated data leads to degeneration into a restricted high-frequency sub-distribution, resulting in a loss of generative diversity and associated capability.
-Argument: We argue that this paradigm’s effectiveness is quite limited, for two reasons: i) Insufficient Capacity.
-Self-enhancement doesn’t increase capacity $C(\mathcal{A})$, negating Premise 1 in Sec.
-2.2.
-This can be regarded as a form of self-training (Feng et al., 2023) minimizing $\text{KL}[\mathcal{A}^{t-1}(\mathbf{x},\mathbf{y})||\mathcal{A}^{t}(\mathbf{x},\mathbf{y})]$, where KL is the Kullback–Leibler divergence and $\mathcal{A}^{t}(\mathbf{x},\mathbf{y})$ is the internal data distribution in $\mathcal{A}$ at iteration $t$, which elicits and reinforces latent capability but cannot expand it, as it is upper bounded by the capacity (Sharma et al., 2024a).
-ii) Mode Collapse might even decrease model capacity (Seddik et al., 2024).
+LABEL:tab:aif_result in Appendix shows that continual training on self-generated data progressively degrades performance as model size increases.
+As a result, the system optimizes for self-preference rather than objective utility, and thus, even if $C(\mathcal{A})$ remains, the capability-capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$ cannot be reduced, or would even increase, undermining Premise 2.
 3.3 Weak-to-Strong Generalization
-Weak-to-Strong Generalization (W2SG) (Burns et al., 2023): This paradigm involves $K$ models, $\mathcal{A}_{1},\!\dots\!,\mathcal{A}_{K}$, with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, to solve task $\mathbf{x}$ that requires superintelligence.
-The goal is to have each $\mathcal{A}_{i}$ generate a solution $\mathbf{y}_{i}$, then fine-tune $\mathcal{A}_{i+1}$ using $\mathbf{y}_{i}$, $i\!=\!1,\!\dots\!,K\!-\!1$ in a progressive manner (called bootstrapping).
-Therefore, each $\mathcal{A}_{i+1}$ surpasses its predecessor $\mathcal{A}_{i}$.
-The key insight is to bridge the internal knowledge $C(\mathcal{A}_{i})$ gained through pretraining with the demanded capability of the target task $U(\mathbf{x})$, eliciting latent knowledge (Mallen et al., 2024) from more advanced models rather than teaching them entirely new skills.
-Below are three key directions:
-(1) Vanilla W2SG (Burns et al., 2023; Sang et al., 2024; Liu & Alahi, 2024; Yang et al., 2024c): The weaker $\mathcal{A}_{i-1}$ with lower capacity generates a solution $\mathbf{y}_{i-1}$ to supervise a stronger $\mathcal{A}_{i}$.
-Performance can be improved by adding an auxiliary confidence loss or incorporating model ensemble.
+Figure 3: Training accuracy of Qwen2.5-7B-Instruct on CosmosQA with different ratios of noisy (flipped labels) training samples.
+LLMs show a stronger tendency to fit noise, except in the range of 0.4 to 0.6 (Green), where labels are heavily randomized.
+In contrast, Fig.
+5 shows that smaller models struggle more with noise.
+More results are in Tab.
+A.9 and Tab.
+LABEL:tab:noise_sciq.
+Weak-to-Strong Generalization (W2SG) (Burns et al., 2023): This paradigm also involves $K$ models but with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, to solve task $\mathbf{x}$ that requires superintelligence.
+The goal is to have each $\mathcal{A}_{k}$ generate a solution $\mathbf{y}_{k}$, then fine-tune $\mathcal{A}_{k+1}$ using $\mathbf{y}_{k}$, $k\!=\!1,\!\dots\!,K\!-\!1$ in a progressive manner (called bootstrapping).
+Therefore, each $\mathcal{A}_{k+1}$ surpasses its predecessor $\mathcal{A}_{k}$.
+The key is to bridge internal knowledge $C(\mathcal{A}_{k})$ with the capability demanded of the target task, eliciting latent knowledge Mallen et al.
+(2024) from more advanced models rather than teaching them entirely new skills.
+Below are three key directions: (1) Vanilla W2SG Burns et al.
+(2023); Sang et al.
+(2024); Liu and Alahi (2024); Yang et al.
+(2024c): The weaker $\mathcal{A}_{i-1}$ with lower capacity generates $\mathbf{y}_{i-1}$ to supervise a stronger $\mathcal{A}_{i}$.
+Performance can be improved with an auxiliary confidence loss or a model ensemble.
 (2) Co-Generalization (Lyu et al., 2024): Both the weaker $\mathcal{A}_{i-1}$ and the stronger $\mathcal{A}_{i}$ jointly produce synthetic solutions $\mathbf{y}$, so that $\mathcal{A}_{i}$’s capacity is fully utilized in providing better and less noisy supervision signals.
 (3) Reward Model Generalization (Sun et al., 2024): Not the policy model $\mathcal{A}$, but the reward model is supervised in the W2SG schema, grounded in the intuition that evaluation is easier than generation (Naor, 1996).
-The supervised reward model then provides signals for the policy model with the same capacity in the subsequent alignment process.
-Figure 3: Training accuracy of Qwen2.5-7B-Instruct on CosmosQA with different ratios (0.0 to 1.0) of noise (flipped labels) in the training set.
-The model shows a stronger tendency to fit noise due to its large parameter size, except in the range of 0.4 to 0.6 (Green), where labels are heavily randomized.
-In contrast, Fig.
-5 shows that smaller models struggle more with noise.
-More results and discussions are presented in Tab.
-A.4 and Tab.
-LABEL:tab:noise_sciq in Appendix.
-Limitations: While W2SG is relatively practical and requires no human intervention, relying on synthetic signals from weaker $\mathcal{A}_{i-1}$ to supervise stronger $\mathcal{A}_{i}$ can lead to unintended behavior, particularly when there is a significant gap between $\mathcal{A}_{i-1}$ and $\mathcal{A}_{i}$.
+The supervised reward model then provides signals for the policy model in the subsequent alignment process.
+Limitations: While W2SG is relatively practical, relying on synthetic signals from weaker $\mathcal{A}_{i-1}$ can lead to unintended behavior, particularly when there is a significant gap $\Delta^{(U)}\!=\!|U(\mathcal{A}_{i})\!-\!U(\mathcal{A}_{i-1})|$.
 This stems from two problems: (1) Imitation of Weak Supervision (Burns et al., 2023).
-The stronger $\mathcal{A}_{i+1}$ may merely imitate how the weaker $\mathcal{A}_{i}$ behave, resulting in a perfect replication of $\mathcal{A}_{i}$ with limited improvement gap (Christiano et al., 2022) instead of generalization.
-Besides, strong $\mathcal{A}_{i+1}$ tends to overfit errors/bias from $\mathcal{A}_{i}$ when synthetic signals contain systematic errors, hurting performance.
+The stronger $\mathcal{A}_{i+1}$ merely imitates or overfits the weaker $\mathcal{A}_{i}$’s behaviors and errors, resulting in a perfect replication of $\mathcal{A}_{i}$ with limited improvement gap Christiano et al.
+(2022) instead of generalization.
 As shown in Fig.
 3, we observe that a highly capable AI is more susceptible to such errors.
-(2) Task Salience in the Strong AI: W2SG is effective only when the task $\mathbf{x}$ being elicited is internally “salient” to $\mathcal{A}_{i+1}$.
-The evidence lies in that in tasks where W2SG succeeds, simple 5-shot prompting also performs comparably (Burns et al., 2023).
+Although Christiano et al.
+(2022) shows that larger $\mathcal{A}_{i}$ agrees less with smaller $\mathcal{A}_{i-1}$’s errors, in superintelligence-level tasks, it’s infeasible to determine whether $\mathcal{A}_{i}$ or $\mathcal{A}_{i-1}$ is correct.
+Tab.
+A.9 demonstrates this: when all synthetic labels are flipped, $\mathcal{A}_{i}$ achieves 100% training accuracy but poor test accuracy.
+This imitation occurs when $\Delta^{(U)}$ is large, invalidating Premise 2.
+(2) Task Salience in the Strong AI: W2SG is effective only when the task $\mathbf{x}$ being elicited is internally “salient" to $\mathcal{A}_{i+1}$.
+The evidence lies in that in tasks where W2SG succeeds, simple 5-shot prompting also performs comparably Burns et al.
+(2023).
 Eliciting non-salient capabilities is more challenging.
-For safety, it’s important that risks caused by current AI may largely differ from those exhibited by future ASI.
-Argument:  We believe that the current W2SG practice is insufficient for achieving superalignment.
-From the limitations discussed above, the key issue is the gap between the weak $\mathcal{A}_{i-1}$ and strong $\mathcal{A}_{i}$, $\delta=|U(\mathcal{A}_{i})-U(\mathcal{A}_{i-1})|$.
-Imitation of weak supervision occurs when $\delta$ is large.
-Although (Christiano et al., 2022) shows that larger $\mathcal{A}_{i}$ agrees less with the errors of smaller $\mathcal{A}_{i-1}$ (or this can be reinforced through confidence loss), in superintelligence-level tasks $\mathbf{x}$, it is infeasible to determine whether $\mathcal{A}_{i}$ or $\mathcal{A}_{i-1}$ is correct ($\mathcal{A}_{i}$’s self-confidence may be biased).
-Tab.
-A.4 in Appendix demonstrates this: when all synthetic labels are flipped, $\mathcal{A}_{i}$ achieves 100% training accuracy but poor test accuracy.
-When the gap $\delta$ is too large, it violates Premise 2 in Sec.
-2.2, causing discussed issues, while a small $\delta$ leads to the Task Salience problem, where the difficult task is not salient for $\mathcal{A}_{i}$, hindering its ability to uncover novel solutions.
-4 A Novel Conceptual Framework
+A small $\delta$ leads to this problem, where the difficult task is not salient for $\mathcal{A}_{i}$, violating Premise 1.
+We can find that existing methods fail to meet both Premises 1 & 2. We argue that solutions exist and detail them in Sec. 4. More analyses of these limitations are provided in Appendix A.6.
+4 Alternating Optimization Framework
+In Sec.
+3, we analyzed that existing paradigms fail to meet Premise 1 & Premise 2 by identifying their limitations.
+Moreover, they focus solely on task competence while overlooking unpredictable risks, especially when sensitive scenarios $\mathbf{x}$ are unknown.
+To address this, we first argue that, beyond existing paradigms, solutions enabling superalignment’s feasibility do exist (Sec.
+4.1), introduce two essential principles critical to satisfying the premises (Sec.
+4.2), and then present ACCO, a conceptual framework that operationalizes these principles (Sec.
+4.3), jointly optimizing competence and conformity (arg.3), as illustrated in Fig.
+4, and explore potential methods of evaluation (Sec.
+4.4).
 Figure 4: Overview of our proposed conceptual framework.
-The superscripts $+$ and -- indicate aligned and unaligned AI, respectively.
+The superscripts ++ and - indicate aligned and unaligned AI, respectively.
 Left: (1) Mere competence enhancement might cause risky AI behavior.
-(2) Directly aligning ASI with capability beyond human score is challenging.
+(2) Directly aligning ASI with capability beyond human scores is challenging.
 (3) A training path by alternating optimization of competence and conformity.
 Right: Three potential paradigms for evaluating the success of superalignment targeting at $\mathbf{x}$ beyond human understanding.
-In Sec.
-3, we analyzed that existing paradigms are inadequate for achieving superalignment due to the limitations identified.
-Moreover, they focus on task competence, i.e., deriving better solutions $\mathbf{y}$ for task $\mathbf{x}$ beyond human capability, while failing to address unpredictable risks (producing safe behavior), as the sensitive scenario $\mathbf{x}$ might be unknown.
-To handle these problems, we discuss the last subargument in Sec.
-2.2 here: superalignment should be achieved via simultaneous optimization of task competence and value conformity, as illustrated in Fig.
-4.
-We first introduce two essential principles for superalignment (Sec.
-4.1), then describe our conceptual framework for superalignment (Sec.
-4.2) and explore potential methods of evaluating success (Sec.
-4.3).
+4.1 Feasibility of Superalignment
+Based on definitions in Sec. 2, superalignment is achievable if Premises 1 & 2 hold. We show these can be readily satisfied beyond existing paradigms and operationalized via the framework in Sec. 4.3.
+Premise 1: Based on the description in Sec.
+3, we believe that superhuman capacity, $C(\mathcal{A}_{\text{ASI}})\!\gg\!C(\mathcal{H})$, can be achieved through model/data scaling.
+The evidence lies in the fact that scaling laws (Kaplan et al., 2020; Gao et al., 2023) still remain effective (OpenAI et al., 2024).
+It has been shown that roughly every seven months, the complexity of tasks AI can perform (measured by human time required to complete them, called time horizon), doubles (Kwa et al., 2025).
+Beyond conventional scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), alternative forms of scaling (Xiao et al., 2024; Wu and Tang, 2024), e.g., inference scaling (Pope et al., 2023; Brown et al., 2024), are being discovered to further boost capacity.
+Despite concerns about data exhaustion, humans continually generate data, especially when assisted by generative AI (Rafiq et al., 2025; Crawl, 2025).
+These facts strongly support Premise 1.
+In fact, once $\mathcal{A}$ masters all human data, which is largely true given the data volume of current AI, $C(\mathcal{A})$ already far exceeds $C(\mathcal{H})$.
+Premise 2: When $C(\mathcal{A}_{\text{ASI}})$ is obtained, a moderate gap $\delta_{\mathcal{A},\mathcal{H}}\!=\!|\mathbb{D}_{U}\!-\!\mathbb{D}_{C}|$ should be closed without golden signals.
+While direct access to golden utility $U(\mathcal{H})$ is infeasible, the capability measure $\mathbb{D}_{U}$ can be approximated via surrogate evaluation (see Sec.
+4.4).
+Thus, minimizing $\delta_{\mathcal{A},\mathcal{H}}$ remains feasible in through a long-term developmental path, if high-quality (semi) synthetic supervision signals can be constructed.This capability expansion mirrors teaching a student multiplication by leveraging addition skills, akin to label propagation (He et al., 2019) and easy-to-hard generalization (Sun et al., 2024), where AI can be trained to solve level-5 math problems using only data on easier (level 1–3) ones (Burns et al., 2023).
 4.2 Two Essential Principles
-As defined in Sec.
-2.1, the superalignment problem can be transformed into constructing supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$.
-Since $\mathcal{S}$ is (partially) synthetic, its quality cannot be guaranteed.
-While there are several limitations in existing paradigms, two key issues stand out: (1) noise in synthetic signals and (2) mode collapse.
+As justified above, the only prerequisite to manifest superalignment is realizable lies in the high-quality supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$.
+The three paradigms in Sec.
+3 fail since they cannot guarantee such an $\mathcal{S}$, but only offer supervision with limited quality (due to human dependency in Sandwiching) and high noise (collapse and self-preference in Self-Enhancement, and weak imitation in W2SG).
 To mitigate them, we propose two guiding principles for constructing $\mathcal{S}$.
 Principle 1: Determining an appropriate capability and capacity gap.
-Our ultimate goal is to develop a solution $\mathbf{y}$ for tasks $\mathbf{x}$ that require superhuman capacity $\tau\!\gg\!C(\mathcal{H})$ (Sec.
-2.2) and capability $U^{*}$ ($\delta\!=\!|U^{*}-U(\mathcal{H})|$ is too large).
-Large $\tau$ and $\delta$ cause several problems.
-For Sandwiching, such a hard $\mathbf{x}$ would be beyond human understanding, and candidate solutions generated by $\mathcal{A}$ with sufficient capacity, $C(\mathcal{A})\!=\!\tau$, be highly deceptive, misleading human judges.
-Our results (Tab.
+As discussed above, we need to develop solutions for tasks that require superhuman capacity $\Delta^{(C)}\!=\!\mathbb{D}_{C}[C(\mathcal{A}_{\text{ASI}})\!-\!C(\mathcal{H})]$ (Sec.
+3) and capability $\delta_{\mathcal{A},\mathcal{H}}\!=\!|\mathbb{D}_{U}[U(\mathcal{A}_{\text{ASI}}),U(\mathcal{H})]\!-\!\Delta^{(C)}|$.
+With $U(\mathcal{H}),U(\mathcal{A})$ fixed, a large $\tau$ (corresponding to a small $\delta_{\mathcal{A},\mathcal{H}}$), indicates an overly hard task, causing noisy synthetic signals, e.g., low-quality ones from Sandwiching, as shown in our results (Tab.
 3.1, Tab.
-A.4, Sec.
-A.2.2, Sec.
-A.3.1) show that capability mismatches between debaters can bias the judge’s decisions in favor of the stronger sneaky debater.
-For W2SG, when $|C(\mathcal{A}_{i+1})\!-\!C(\mathcal{A}_{i})|$ is large, imitation of weak supervision occurs.
-As reported in (Burns et al., 2023), Fig.
+A.9, Sec.
+A.7.2, Sec.
+A.8.1), where the capability mismatches between debaters can mislead the judge’s decisions in favor of the stronger sneaky debater.
+Besides, when $\Delta^{(C)}$ is large, imitation of weak supervision occurs, which makes the student $\mathcal{A}_{i+1}$’s performance lag (negating Premise 2), as reported in Burns et al.
+(2023), Fig.
 5, Tab.
-A.4 and Tab.
-LABEL:tab:noise_sciq, such a large gap makes the student $\mathcal{A}_{i+1}$’s performance lag (that’s why we assume Premise 2 in Sec.
-2.2).
-However, a too small gap is also not optimal.
-Since the initial capacity $C(\mathcal{A}_{1})$ falls far below $\tau$, making the target task $\mathbf{x}$ non-salient in the student models $\mathcal{A}_{1},\mathcal{A}_{2},\dots$, as mentioned in Sec.
-3.3 (in this case, $|U^{*}\!-\!U(\mathcal{A}_{i})|\!>\!|C(\mathcal{A}_{i+1})\!-\!C(\mathcal{A}_{i})|$), failing to elicit necessary capability.
+A.9 and Tab.
+LABEL:tab:noise_sciq.
+However, a $\Delta^{(C)}$ too small is also suboptimal, as the initial capacity $C(\mathcal{A}_{1})$ falls far below $\Delta^{(C)}$, making the target task non-salient in the student models $\mathcal{A}_{1},\mathcal{A}_{2},\dots$, as mentioned in Sec.
+3.3, failing to elicit necessary capabilities.
 Therefore, an appropriate gap is needed.
 Principle 2: Increasing the diversity of signals.
-Synthetic data $\mathcal{S}$ would cause significant mode collapse (Shumailov et al., 2024; Gerstgrasser et al., 2024) and self-preference (Wataoka et al., 2024) and bias (Ziems et al., 2024) (for self-enhancement and W2SG), which not only fails to expand the model’s capability but may also degrade it.
-Instead of relying on one single data source (human or AI), we propose to combine multiple sources to construct the signals, that is, $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}^{k}\}\bigcup\{\mathbf{x}_{i},\mathbf{y}_{i}^{j}\}$ where $\mathbf{y}_{i}^{k}\!\sim\!\mathcal{A}_{k}(\mathbf{x}_{i})$ and $\mathbf{y}_{i}^{j}\!\sim\!\mathcal{H}_{j}(\mathbf{x}_{i})$.
+While not as a standalone solution to value alignment [^6] We acknowledge the extensive literature on value alignment (Anwar et al., 2024; Sorensen et al., 2024; Hadfield-Menell et al., 2016, 2019; Fisac et al., 2020; Barez and Torr, 2023).
+In contrast, our work formalizes and investigates alignment for superhuman systems where ground truth supervision is infeasible., diversity of supervision signals plays an important role.
+As analyzed above, the low-quality $\mathcal{S}$ also originates from model collapse (Shumailov et al., 2024; Gerstgrasser et al., 2024), self-preference (Wataoka et al., 2024) and bias (Ziems et al., 2024), even if we find an appropriate gap $\tau$, which not only fails to expand the model’s capability but may also degrade it.
+Instead of relying on single data source (human or AI), we propose to combine multiple sources to construct the signals, that is, $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}^{k}\}\bigcup\{\mathbf{x}_{i},\mathbf{y}_{i}^{j}\}$ where $\mathbf{y}_{i}^{k}\!\sim\!\mathcal{A}_{k}(\mathbf{x}_{i})$ and $\mathbf{y}_{i}^{j}\!\sim\!\mathcal{H}_{j}(\mathbf{x}_{i})$.
 Diverse signal sources (e.g., heterogeneous models or humans with distinct backgrounds) could help mitigate these issues (Du et al., 2024; Chen et al., 2024).
@@
 LABEL:fig:rewardbench_evaluation_all, Sec.
-A.2.4 and Sec.
-A.3.3 demonstrate that noise of synthetic data negatively affects performance while multiple sources would help.
-Diversity can be achieved by (1) Human-in-the-Loop Feedback (Hejna III & Sadigh, 2023; Retzlaff et al., 2024), (2) Alternative Backbone Models (Kim et al., 2023), and (3) Role-Playing (Shao et al., 2023; Wang et al., 2023b)
-4.2 Our Framework
-Following the principles above, we illustrate a conceptual framework for superalignment, as shown in Fig. 4 (left).
-Define $\mathbf{x}^{*}$ as the hard task (e.g., prove the Riemann Hypothesis), and $\mathcal{A}_{1},\!\dots\!,\mathcal{A}_{K}$ as $K$ models with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, as those used in W2SG (Sec.
+A.7.4 and Sec.
+A.8.3 demonstrate that multiple sources would reduce noise and avoid bias or collapse.
+4.3 The ACCO Framework
+To operationalize the two principles above, we propose a conceptual Alternating Competence and Conformity Optimization (ACCO) Framework, as shown in Fig. 4 (left).
+Define $\mathbf{x}^{*}$ as the hard task (e.g., prove the Riemann Hypothesis), and $\{\mathcal{A}_{k}\}_{i=1}^{K}$ as $K$ models with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, as those used in W2SG (Sec.
 3.3).
-The capability gap is defined as $\delta_{i}\!=\!|C(\mathcal{A}_{i+1})-C(\mathcal{A}_{i})|$.
-Rather than directly requiring $\mathcal{A}_{i}$ to solve $\mathbf{x}^{*}$, we can find or construct a series of predecessor tasks, $\mathbf{x}_{1},\dots,\mathbf{x}_{K-1}$ [^3] Each task may not be a single instance but rather a class of tasks encompassing multiple instances.
-For clarity, we describe it using a single $\mathbf{x}$ as a representative.
+Rather than directly requiring $\mathcal{A}_{i}$ to solve $\mathbf{x}^{*}$, we can find or construct a series of predecessor tasks, $\{\mathbf{x}_{i}\}_{i=1}^{K-1}$ [^7] Each may not be a single but a class of tasks with multiple instances.
+For clarity, we use a single $\mathbf{x}$.
 in the same domain (like mathematics) with increasing difficulty, e.g., $\mathbf{x}_{1}$ is a basic arithmetic question, $\mathbf{x}_{2}$ is a calculus question, $\mathbf{x}_{K\!-\!1}\!=\!$ “Prove Mordell conjecture”.
-The difficulty gap between $\mathbf{x}_{i}$ and $\mathbf{x}_{i+1}$ should match the capacity gap $\delta_{i}$, so that task $\mathbf{x}_{i+1}$ is salient in $\mathcal{A}_{i+1}$.
-Then we propose to conduct a progressive training process, highlighting two key concepts: Gap Annealing and alternating competence enhancement and value alignment.
-The complete process is described as below:
+The difficulty gap between $\mathbf{x}_{i}$ and $\mathbf{x}_{i+1}$ should match the capacity gap $\Delta^{(C)}_{i}$, so that $\mathbf{x}_{i+1}$ is salient in $\mathcal{A}_{i+1}$.
+Then we propose a progressive training process, highlighting two keys: Gap Annealing and alternating competence enhancement and value alignment [^8] Our framework is value-type agnostic, as the utility function $U(\mathbf{x})$ can reflect either universal (e.g., fairness, non-discrimination) or culturally grounded values, depending on the supervision signal.
+This flexibility supports ongoing research on pluralistic alignment Alkhamissi et al.
+(2024); Anwar et al.
+(2024)..
+The complete process is as below:
 For $i=1,\dots,K\!-\!1$, perform the following steps:
-Step 1: Supervise $\mathcal{A}_{i}$ by $\mathcal{S}_{i}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}$, where $\mathbf{y}_{i}$ could be either a ground truth or synthetic one produced from the previous iteration.
-As long as $\mathcal{A}_{i+1}$ has not exceeded the human intelligence bound, ground truth remains available.
-Step 2: Without safety constraints, $\mathcal{A}_{i+1}$ with higher capability may exhibit harmful behaviors, either proactively (Bengio et al., 2024b) or passively (Chao et al., 2023), defined as $\mathcal{A}_{i+1}^{-}$.
+Step 1: Supervise $\mathcal{A}_{i}$ by $\mathcal{S}_{i}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}$, where $\mathbf{y}_{i}$ is either a ground truth or synthetic one produced from the previous iteration.
+When $\mathcal{A}_{i+1}$ is below human intelligence, ground truth remains available.
+Step 2: Without value alignment, $\mathcal{A}_{i+1}$ with higher capability may exhibit harmful behaviors, either proactively (Bengio et al., 2024b) or passively (Chao et al., 2023), defined as $\mathcal{A}_{i+1}^{-}$.
 Deploy $\mathcal{A}_{i+1}^{-}$ in real or experimental environments, where its potential risks are identified through observation or red-teaming (Perez et al., 2022), uncovering hazardous scenarios $\hat{\mathbf{x}}$.
-Derive the corresponding safe behaviors $\hat{\mathbf{y}}$ using $\mathcal{A}_{1},\dots,\mathcal{A}_{i}$ and human participants $\mathcal{H}$.
+Derive safe behaviors $\hat{\mathbf{y}}$ with $\{\mathcal{A}_{k}\}_{k=1}^{i}$ and human $\mathcal{H}$.
 Step 3: Supervise $\mathcal{A}_{i+1}^{-}$ via $\{\hat{\mathbf{x}},\hat{\mathbf{y}}\}$ to obtain the value-aligned $\mathcal{A}_{i+1}^{+}$. $\mathcal{A}_{i+1}\leftarrow\mathcal{A}_{i+1}^{+}$.
-Step 4: Adaptively determine (increase or decrease) the capacity gap $\delta_{i+1}$ according to the real capacity $C(\mathcal{A}_{i+1})$ and utility $U(\mathcal{A}_{i+1})$ evaluated.
+Step 4: Adaptively determine (increase or decrease) the capacity gap $\Delta_{i+1}^{(C)}$ according to the real capacity $C(\mathcal{A}_{i+1})$ and utility $U(\mathcal{A}_{i+1})$ evaluated.
 This framework offers two advantages.
 (1) Dynamic and flexible capacity gap.
-$\delta$ can be adjusted using an annealing process (Fu et al., 2019; Huang et al., 2020).
-In the early stages, a smaller $\delta$ helps prevent imitation of weak supervision (when noise is high).
+$\Delta^{(C)}$ can be adjusted using an annealing process (Fu et al., 2019; Huang et al., 2020), and determined by search-based optimization (e.g., Monte Carlo tree search (undefinedwiechowski et al., 2022)) to adaptively explore alignment strategies (Khanov et al., 2024).
+In the early stages, a smaller $\Delta^{(C)}$ helps prevent imitation of weak supervision (when noise is high).
 As the process advances, the gap can be gradually increased to ensure task saliency, assuming $U(\mathcal{A}_{i})$ is sufficiently high and data noise decreases.
-(2) Exposure to unknown hazardous scenarios.
-As AI surpasses human capabilities, its potentially harmful behaviors and triggers become unpredictable.
-Rather than developing a full ASI with extreme risks (Critch & Krueger, 2020; Goldstein et al., 2023) and aligning it afterward, we can expose risk in a controlled environment, where (a) models are not yet capable of causing catastrophic harm, and (b) researchers have time and ability to mitigate them (Cotra, 2021).
-Note that diversity enhancement approaches can be adopted in each iteration, fostering exploration and safety.
+(2) Exposure to unknown hazardous scenarios (to support arg.3).
+As AI surpasses human capabilities $U(\mathcal{A})\gg U(\mathcal{H})$, its potentially harmful behaviors and triggers become unpredictable.
+Rather than developing a full ASI with extreme risks (Critch and Krueger, 2020; Goldstein et al., 2023) and aligning it afterward, we can expose risk in a controlled environment, gradually increasing its utility $U(\mathcal{A}_{1})\!<\!\cdots\!<\!U(\mathcal{A}_{i+1})$ where (a) models are not yet capable of causing catastrophic harm, and (b) researchers have the time and ability to mitigate them (Cotra, 2021), providing time to correct misaligned behaviors before the model becomes powerful enough to cause catastrophic scenarios $\hat{\mathbf{x}}$ (e.g., a treacherous turn).
+In each iteration, diversity enhancement approaches can be adopted in each iteration, fostering exploration and safety.
+This can be achieved by (1) collaborative and competitive interactions among AI (Khan et al., 2024) to uncover failure cases and solutions, (2) human-in-loop feedback (Hejna III and Sadigh, 2023; Retzlaff et al., 2024), (3) alternative backbone models (Kim et al., 2023), or (4) role-playing (Shao et al., 2023; Wang et al., 2023b).
+More details are presented in Appendix A.4
 4.4 Superalignment Evaluation
-The evaluation of superalignment is challenging, as the ground truth $\mathbf{y}$ is unavailable and even unassessable. We discuss three potential paradigms, as shown in Fig. 4 (right).
-Surrogate Evaluation:  Estimating AI’s performance in real-world cases by that evaluated in simplified cases (Ruan et al., 2024).
-By allowing AI to operate within these controlled environments, we may examine AI behavior and identify edge cases.
-For example, an autonomous navigation AI can be tested in a virtual environment simulating diverse traffic, weather, and pedestrian scenarios (Dosovitskiy et al., 2017; Kiran et al., 2021).
-Extrapolation:  Uncovering emergent behaviors and potential misalignments as AI scales (McKenzie et al., 2023).
-This method continuously evaluates the model as it scales to detect performance shifts.
+The evaluation of superalignment is challenging, as the ground truth $\mathbf{y}$ is unavailable and unassessable.
+We discuss three potential paradigms, as shown in Fig.
+4 (right), which use approximated utility $\hat{U}(\mathbf{y})$ as a proxy for the golden utility $U(\mathbf{y})$ to obtain an empirical estimate of $\mathbb{D}_{U}$.
+(1) Surrogate Evaluation:  Estimating AI’s performance in real-world cases by evaluation in simplified cases (Ruan et al., 2024).
+By allowing AI to operate within controlled environments, we may examine AI behavior.
+For example, an autonomous navigation AI can be tested in a virtual environment simulating diverse traffic and pedestrian scenarios Dosovitskiy et al.
+(2017); Kiran et al.
+(2021).
+(2) Extrapolation:  Uncovering emergent behaviors and potential misalignments as AI scales McKenzie et al.
+(2023).
+This method continuously evaluates AI as it scales.
 Perez et al.
-(2023) propose using AI to generate evaluation datasets, identifying unexpected behaviors.
-Additionally, we suggest tracking alignment failures when scaling, as misalignments may follow a U-shaped pattern, emerging after initial peaks or valleys in tuning.
-Peer Review:  Domain experts can qualitatively assess model outputs and reasoning for alignment.
-Peer review, shown to enhance academic performance (Double et al., 2020), has also been used to mitigate noisy supervision (Lu & Kong, 2023).
-Similarly, expert panels can evaluate a paper draft on technical accuracy and clarity.
-Engaging multiple reviewers ensures a more objective assessment of a model’s performance and alignment with human values.
-4.4 Potential Implementation
-The conceptual framework described in Sec.
-4.2 is promising, but challenges remain in implementing it in practice.
-In this section, we briefly outline some potential directions.
-For more detail, please refer to Sec.
-A.1.
-Enhancing Data Diversity  Data diversity is essential to avoid mode collapse (Principle 2).
-Collaboration among AI models with different specialties improves supervision (Guo et al., 2024), while competition exposes alignment weaknesses and solutions (Khan et al., 2024).
-However, filtering is essential to prevent undesirable behaviors and avoid noisy signals, as discussed in Sec.
+Perez et al.
+(2023) propose using AI to generate evaluation datasets and identify unexpected behaviors.
+We also suggest tracking alignment failures when scaling, as misalignments may follow a U-shaped pattern.
+(3) Peer Review:  Domain experts can qualitatively assess model outputs for alignment.
+Peer review, shown to enhance academic performance Double et al.
+(2020), has also been used to mitigate noisy supervision Lu and Kong (2023).
+This is similar to experts who evaluate a paper draft on technical accuracy and clarity.
+See Appendix.
+A.5 for more detailed and formal descriptions.
+5 Alternative Views
+While we argue that research on superalignment should begin now (Sec.
+2.2), several alternative views persist.
+We summarize key points below and address them in Appendix A.1.
+Some claim that superalignment is unattainable, arguing that without informed oversight, supervising more-capable AI systems is infeasible Burns et al.
+(2023).
+Others view superalignment as an emergent property of model scaling Kaplan et al.
+(2020); Wei et al.
+(2023), yet this overlooks risks like deceptive behavior Carranza et al.
+(2023) and inverse scaling McKenzie et al.
+(2023).
+A third concern is that superalignment is hard to empirically verify, due to the lack of ground truth Yin et al.
+(2024).
+We address this by proposing potential evaluation methods in Sec.
+4.4.
+Some argue there is no fundamental difference from current alignment, viewing superalignment as a continuation of standard alignment Fisac et al.
+(2020); however, it has distinct challenges under insufficient human oversight, detailed in Appendix A.3.
+Finally, others suggest superalignment can be achieved through inference scaling Muennighoff et al.
+(2025), though this cannot expand capacity $C(\mathcal{A})$ (Premise 1 in Sec.
+3).
+6 Conclusion
+In this paper, we argue that superalignment is achievable and research on it should advance immediately, through alternating optimization of competence and conformity.
+We provide its formal definition rooted in scalable oversight and the capability–capacity gap, and review existing paradigms to expose their limitations.
+To address these, we propose two novel principles.
+(1) Determining an appropriate gap, and (2) increasing the diversity of signals, and then outline the ACCO framework, with a potential evaluation method to align ASI.
+By laying this foundation, we aim to illuminate the landscape for addressing the critical challenge of aligning future ASI with human values.
+Appendix A Appendix
+A.1 Alternative Views
+Superalignment is Unattainable  Absence of ASI and the lack of informed oversight have led some to argue that superalignment is unattainable due to $C(\mathcal{A})\gg C(\mathcal{H})$ (Sec.
+2).
+Given the current limitations, humans cannot reliably supervise AI systems $\mathcal{A}$ when their capabilities exceed human expertise, making it difficult to prevent $\mathcal{A}_{\text{ASI}}$ from deceiving its weaker teacher [31].
+Moreover, while existing alignment methods [204, 148, 71, 196, 66] offer limited success in aligning $\mathcal{A}_{\text{ANI}}$, they face scalability and reliability issues.
+However, we show in Sec.
+4.4 (Superalignment Evaluation) that $\mathbb{D}_{U}$ can be approximated without golden label.
+Superalignment is an Emergent Property  Building on scaling laws [93, 186, 128], one could argue that superalignment might emerge as a byproduct of intelligence expansion rather than requiring proactive intervention.
+However, concerns about deceptive alignment [33, 75] and inverse scaling [117] persist.
+Formally, increased capacity $C(\mathcal{A})$ does not guarantee aligned capability $U(\mathcal{A})$, and superalignment explicitly requires minimizing the capability-capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$.
+As models develop advanced capability, new challenges arise [174, 32, 68].
+Relying solely on emergent properties poses risks, including unpredictability in AI behavior.
+[…]
+No Fundamental Difference from Current Alignment  Superalignment can simply be seen as advanced alignment, relying on current feedback techniques [74, 58].
+We argue instead that superalignment arises when $C(A)\gg C(H)$ and humans can no longer supervise effectively.
+Unlike conventional alignment which minimize capability gap $|U(\mathcal{A})-U(\mathcal{H})|$, our framework addresses the divergence between capability-capacity $|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})-\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})|$, requiring long-term, cross-generational methods beyond typical fine-tuning.
+For more details about the difference between superalignment and standard alignment, please refer to Appendix A.3.
+Superalignment Can Be Achieved through Inference Scaling  Some argue that superalignment may emerge from inference-time scaling rather than model scaling [127, 201].
+While inference scaling shows promise on structured tasks [69], and may improve performance $U(\mathcal{A})$, it does not increase model capacity or internalized knowledge $C(\mathcal{A})$, thus violating Premise 1.
+As a result, the capability–capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$, under the assumption $C(\mathcal{A})\gg C(\mathcal{H})$, cannot be minimized.
+Our Counterargument  As discussed in Sec.
+2.2, we state our position that superalignment is achievable and research on it should advance immediately.
+Although current paradigms have limitations, delaying research would pose more significant risks [54, 79, 21, 67].
+Instead, we propose a simultaneous optimization approach to achieve superalignment, integrating task competence [125] and value conformity [7, 174, 131].
+A.2 Impact Statement
+This work introduces a formal definition of superalignment, connecting existing concepts by exploring their interconnections and limitations.
+We then propose a potential pathway to achieve superalignment, grounded in two fundamental principles: (1) determining an appropriate capability-capacity gap and (2) increasing the diversity of signals.
+Although our work proposes a potential pathway to address key challenges, it still contains three main challenges that should be addressed in the future.
+First, our position relies on assumptions about the speculative nature of ASI, which has not yet been realized.
+This introduces uncertainty regarding the applicability of our method in future scenarios, but we argue its relevance for work that approaches ASI by dominating human performance across multiple criteria and contexts.
+Second, our evaluations are limited in terms of model size and scalability.
+While we demonstrate promising results that address current limitations and highlight the potential of our approach, our experiments are constrained by the use of a 7B parameter model as the largest one tested.
+Lastly, as shown in Fig.
+1, it is important to consider both aligned and unaligned $\mathcal{A}_{\text{ASI}}$ but also cases where ASI is partially aligned with specific objectives—for example, an ASI model aligned with ethics but unaligned with safety.
+Active discussion is needed to deepen the collective understanding of ASI alignment.
+In terms of ethical concerns, (1) Amplified Biases.
+W2SG may propagate biases inherent in training data or weak models.
+(2) Increased Safety Risks.
+More capable models are more susceptible to noise, which may trigger unintended harmful behaviors.
+(3) Other Impacts AI’s biases and potential sycophancy can produce outputs overly aligned with unreliable human feedback, causing social harm.
+We hope that this paper will influence organizations considering the development of superalignment AI models by not only clarifying existing alignment methods and their limitations but also presenting a potential pathway for achieving superalignment.
+By addressing current challenges and proposing a potential solution, we aim to contribute to the broader goal of developing AI models that provide benefits to humanity while minimizing potential harms.
+A.3 Distinguishing Superalignment from Existing Alignment
+In this section, we clarify the core differences between superalignment and existing alignment, especially in terms of their assumptions, goals, and methodological constraints.
+Definition and Objective Existing alignment aims to align an AI system’s behavior with human values or preferences i.e., $U(\mathcal{A})\approx U(\mathcal{H})$.
+This assumes the model has sufficient capacity such that $C(\mathcal{A})\geq C(\mathcal{H})$ and a utility that is comparable to or lower than that of humans such that $U(\mathcal{A})<U(\mathcal{H})$.
+As formalized in Def.
+2.1 in Sec.
+2.1, the objective is to minimize the utility gap, defined by $\mathcal{A}^{*}=\mathop{\text{argmin}}\nolimits_{\mathcal{A}}|U(\mathcal{A})-U(\mathcal{H})|$.
+In contrast, superalignment aims to align an AI system whose utility $U(\mathcal{A})$ is required to significantly exceed that of any human, i.e., $U(\mathcal{A}_{\text{ASI}})\gg U(\mathcal{H})$.
+In such cases, human capability $U(\mathcal{H})$ is no longer a valid alignment reference.
+Instead, the superalignment objective focuses on minimizing the gap between a model’s measured capability and capacity, quantified by $\mathbb{D}_{U}$ and $\mathbb{D}_{C}$, respectively.
+As introduced in Def.
+2.1, this is formalized as $\mathcal{A}^{*}=\mathop{\text{argmin}}\nolimits_{\mathcal{A}}\left|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})]-\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})]\right|$ without changing $C(\mathcal{A})$.
+Key Differences The core difference lies in the supervision approach.
+Existing alignment assumes that humans can provide or verify supervision signals, making techniques such as fine-tuning or reinforcement learning from human feedback (RLHF) feasible.
+Superalignment, however, arises in contexts where human supervision fails; e.g., when evaluating a model’s proof of the Riemann Hypothesis, even expert humans may be unable to assess the correctness of the solution.
+Consequently, existing alignment focuses on aligning models to human-level utility, $|U(\mathcal{A})-U(\mathcal{H})|<\epsilon$, while superalignment aims to align models to their own capacity in a way that conforms to intended utility structures, $|\mathbb{D}_{U}[U(\mathcal{A}),U(\mathcal{H})]-\mathbb{D}_{C}[C(\mathcal{A}),C(\mathcal{H})]|<\epsilon$.
+The two also differ in terms of risk.
+Existing alignment assumes that the human-estimated utility $\hat{U}_{\mathcal{H}}(\mathcal{A})$ is a reliable proxy for the golden utility $U(\mathcal{A})$, and that misalignment primarily results from human misjudgment or bias, with the deviation $|\hat{U}_{\mathcal{H}}(\mathcal{A})-U(\mathcal{A})|$ remaining sufficiently small.
+In contrast, superalignment assumes in domains where $\hat{U}_{\mathcal{H}}(\mathcal{A})$ is undefined or unreliable, such that $|\hat{U}_{\mathcal{H}}(\mathcal{A})-U(\mathcal{A})|$ becomes large due to emergent capabilities that are unpredictable or unverifiable by humans (e.g., treacherous turn).
+Evaluation In terms of evaluation, while existing alignment assumes access to a ground truth $\mathbf{y}$, superalignment lacks such ground truth, which is often unavailable or unassessable, making its evaluation particularly challenging.
+To lay the groundwork for evaluating superalignment, we propose several approximation methods in Sec.
+4.4.
+A.4 Potential Directions in Detail
+The conceptual framework described in Sec. 4.3 is promising, but challenges remain in implementing it in practice. In this section, we provide detailed potential directions.
+Fostering Creativity Through Data Diversity
+Data diversity is essential to avoid mode collapse (Principle 2).
+Formally, let the supervision signal be denoted as $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}^{j}\}$, where $\mathbf{y}_{i}^{j}\sim\mathcal{A}_{j}(\mathbf{x}_{i})$ for models $\mathcal{A}_{j}$ with varying capacities or perspectives, or $\mathbf{y}_{i}^{j}\sim\mathcal{H}_{j}(\mathbf{x}_{i})$ for human agents with diverse backgrounds.
+Diversity in $\mathcal{S}$ helps reduce the bias in $\hat{U}(\mathcal{A})=\mathbb{E}_{\mathbf{x}\sim\mathcal{X}}[\hat{U}(\mathcal{A}(\mathbf{x}))]$.
+Collaboration among heterogeneous AI systems [72] and competition among models [95] expose alignment weaknesses and solutions.
+However, filtering is essential to prevent undesirable behaviors and avoid noisy signals that may increase $|\mathbb{D}_{U}-\mathbb{D}_{C}|$, as discussed in Sec.
 3.1 and Sec.
-3.1.
-Iterative Training Between Teacher and Student  In iterative teacher-student training when $\mathcal{C}(\mathcal{A}_{i-1})<\mathcal{C}(\mathcal{A}_{i})$, a strong teacher $\mathcal{A}^{t-1}_{i}$ generates a solution $\mathbf{y}^{t-1}_{i}$ to supervise a weaker student $\mathcal{A}^{t-1}_{i-1}$, and the supervised weaker student $\mathcal{A}^{t}_{i-1}$ then provides feedback $\mathbf{y}^{t}_{i-1}$ to improve the teacher $\mathcal{A}^{t}_{i}$ (Yuan et al., 2021).
+3.3.
+Iterative Training Between Teacher and Student
+In iterative teacher-student training when $\mathcal{C}(\mathcal{A}_{i-1})<\mathcal{C}(\mathcal{A}_{i})$, a strong teacher $\mathcal{A}^{t-1}_{i}$ generates a solution $\mathbf{y}^{t-1}_{i}$ to supervise a weaker student $\mathcal{A}^{t-1}_{i-1}$, and the supervised weaker student $\mathcal{A}^{t}_{i-1}$ then provides feedback $\mathbf{y}^{t}_{i-1}$ to improve the teacher $\mathcal{A}^{t}_{i}$ [203].
 This cyclical process can enhance alignment but may face challenges from limited diverse feedback (Sec.
@@
 3.2).
-Search-Based Methods Toward Optimal Learning Direction  Search-based methods optimize alignment strategies by exploring possible alignment strategies (Khanov et al., 2024).
-Techniques like Monte Carlo tree search (undefinedwiechowski et al., 2022) or gradient-free optimization (Larson et al., 2019) can guide promising directions.
+Search-Based Method Toward Optimal Learning Direction
+Search-based methods optimize alignment strategies by exploring possible alignment strategies [96].
+Given a fixed model capacity $C(\mathcal{A})$, these methods aim to optimize the realized capability $U(\mathcal{A})$ by searching over policy or alignment configurations to minimize the gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$.
+Techniques such as Monte Carlo tree search [180] or gradient-free optimization [104] can guide promising directions.
 However, these methods may require high computational resources.
-5 Alternative Views
-Superalignment is Unattainable  Absence of ASI and the lack of informed oversight have led some to argue that superalignment is unattainable (Sec.1).
-Given the current limitations, humans cannot reliably supervise AI systems when their capabilities exceed human expertise, making it difficult to prevent $\mathcal{A}_{\text{ASI}}$ from deceiving its weaker teacher (Burns et al., 2023).
-Moreover, while existing alignment methods (Zhou et al., 2023; Rafailov et al., 2023; Gulcehre et al., 2023; Xu et al., 2023; Gou et al., 2024) offer limited success in aligning $\mathcal{A}_{\text{ANI}}$, they face scalability and reliability issues.
-Superalignment is an Emergent Property  Building on scaling laws (Kaplan et al., 2020; Wei et al., 2023; Nam et al., 2024), one could argue that superalignment might emerge as a byproduct of intelligence expansion rather than requiring proactive intervention.
-However, concerns about deceptive alignment (Carranza et al., 2023; Hagendorff, 2024) and inverse scaling (McKenzie et al., 2023) persist.
-As models develop advanced capability, new challenges arise (Taylor et al., 2016; Carlsmith, 2022; Greenblatt et al., 2024b).
-Relying solely on emergent properties poses risks, including unpredictability in AI behavior.
-Our Counterargument  As discussed in Sec.
-2.2, we state our position that superalignment is achievable and research on it should advance immediately.
-Although current paradigms have limitations, delaying research would pose more significant risks (Federspiel et al., 2023; Hendrycks et al., 2023; Bengio et al., 2024a; Greenblatt et al., 2024a).
-Instead, we propose a simultaneous optimization approach to achieve superalignment, integrating task competence (Motwani et al., 2024) and value conformity (Amodei et al., 2016; Taylor et al., 2016; Olah, 2016).
-6 Conclusion
-In this paper, we argue that superalignment is achievable and research on it should advance immediately, through simultaneous and alternating optimization of task competence and value conformity.
-We provide a formal definition of superalignment, originating from scalable oversight.
-Then, we review three existing paradigms and explore their limitations.
-To address the limitation, we propose two fundamental principles as a potential path to superalignment.
-First, determining an appropriate capability and capacity gap.
-Second, increasing the diversity of signals.
-We then outline a conceptual framework for superalignment, including a potential evaluation method to align ASI.
-By laying this foundation, we aim to illuminate the landscape for addressing the critical challenge of aligning future ASI with human values, ensuring its benefits for humanity.
-Impact Statement
-This work introduces a formal definition of superalignment, connecting existing concepts by exploring their interconnections and limitations.
-We then propose a potential pathway to achieve superalignment, grounded in two fundamental principles: (1) determining an appropriate capability and capacity gap and (2) increasing the diversity of signals.
-Although our work proposes a potential pathway to address key challenges, it still contains three main challenges that should be addressed in the future.
-First, our position relies on assumptions about the speculative nature of ASI, which has not yet been realized.
-This introduces uncertainty regarding the applicability of our methods in future scenarios.
-Second, our evaluations are limited in terms of model size scalability.
-While we have demonstrated promising results that address current limitations and highlight the potential of our approach, our experiments are constrained by the use of a 7B parameter model as the largest tested model.
-Lastly, as shown in Fig.
-1, it is important to consider not only $ASI_{unaligned}$ and $ASI_{aligned}$ but also cases where ASI is partially aligned with specific objectives—for example, an ASI model that is aligned with ethics but unaligned with safety.
-Active discussion is needed to deepen our understanding of ASI alignment.
-This paper could influence organizations considering the development of superalignment AI models by not only clarifying existing alignment methods and their limitations but also presenting a potential pathway to achieving superalignment.
-By addressing current challenges and proposing a potential solution, we aim to contribute to the broader goal of developing AI models that provide benefits to humanity while minimizing potential harm.
-Appendix A Appendix
-A.1 Potential Directions in Detail
-Fostering Creativity Through Data Diversity
-Encouraging data diversity can play a pivotal role in fostering creativity and robustness in alignment strategies.
-Multi-AI system collaboration, where AI systems with different specializations or perspectives interact, has shown promise in enabling emergent behaviors and problem-solving approaches beyond the capabilities of individual AI systems (Guo et al., 2024).
-Competition among agents can also drive improvement, as adversarial settings often reveal diverse argument data that finds alignment weaknesses and suggests novel pathways for improvement (Khan et al., 2024).
-However, diversity-based approaches require careful curation and filtering to ensure that resulting behaviors are beneficial and do not inadvertently reinforce undesirable outcomes.
-Iterative Training Between Teacher and Student
-Iterative teacher-student training involves a cyclical process where a more advanced teacher guides a less capable student, and the teacher subsequently obtains feedback from the student for adjustments (Yuan et al., 2021).
-This process can be extended such that the student becomes the teacher in the next iteration, guiding the previous teacher to improve its alignment capabilities.
-This approach enables incremental learning, where each output generated by the AI system builds upon and enhances the alignment capabilities of its predecessor.
-While this method offers scalability and the potential to bootstrap stronger AI systems from weaker supervision signals, challenges such as those posed by W2SG (Sec.
-3.3) and self-enhancement (Sec.
-3.2) may arise due to limited diversity.
-Effective strategies must be devised to mitigate these risks.
-Search-Based Method Toward Optimal Learning Direction
-Search-based methods use optimization techniques to explore the space of possible alignment strategies, seeking an optimal pathway to systematically navigate complex alignment challenges (Khanov et al., 2024).
-These methods are particularly useful when the optimal alignment direction is unclear, as they allow for exploration and experimentation in uncertain spaces.
-Techniques like Monte Carlo tree search (undefinedwiechowski et al., 2022) or gradient-free optimization (Larson et al., 2019) can guide this process by prioritizing regions of the search space with the highest potential for success.
-However, these methods require significant computational resources.
-A.2 Experiment Setup
-A.2.1 General
+A.5 Superalignment Evaluation in Detail
+[…]
+[…]
+Extrapolation:  Uncovering emergent behaviors and potential misalignments as AI scales [117].
+This method continuously evaluates the model as it scales to detect performance shifts.
+Perez et al.
+[141] propose using AI to generate evaluation datasets, identifying unexpected behaviors.
+Additionally, we suggest tracking alignment failures when scaling, as misalignments may follow a U-shaped pattern, emerging after initial peaks or valleys in tuning.
+Another extrapolation-based method may involve training AI models on easy tasks and evaluating their generalization to harder problems [173].
+Formally, let $\{\mathcal{A}_{1},\dots,\mathcal{A}_{K}\}$ be $K$ models with increasing capacities, i.e., $C(\mathcal{A}_{1})<\cdots<C(\mathcal{A}_{K})$.
+Let $\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$ be a corresponding sequence of tasks with increasing complexity.
+For each model-task pair $(\mathcal{A}_{i},\mathbf{x}_{i})$, the capability can be approximated as $\hat{U}(\mathbf{y}_{i})$.
+Extrapolation examines whether $\{\hat{U}(\mathbf{y}_{i})\}_{i=1}^{K}$ exhibits a consistent improvement or U-shaped pattern, revealing potential misalignment as models scale.
+Peer Review:  Domain experts can qualitatively assess model outputs and reasoning for alignment.
+Peer review, shown to enhance academic performance [48], has also been used to mitigate noisy supervision [112].
+Similarly, expert panels can evaluate a paper draft on technical accuracy and clarity.
+Engaging multiple reviewers ensures a more objective assessment of a model’s performance and alignment with human values.
+Formally, let $\{\mathcal{H}_{1},\dots,\mathcal{H}_{J}\}$ be $J$ reviewers with their own evaluation function $\hat{U}_{\mathcal{H}_{i}}$.
+For input $\mathbf{x}$, $\mathbf{y}=\mathcal{A}(\mathbf{x})$ is evaluated by all reviewer $\hat{U}(\mathbf{y})=\frac{1}{J}\sum_{i=1}^{J}\hat{U}_{\mathcal{H}_{i}}(\mathbf{y})$.
+A.6 A Comprehensive Overview of Limitations in Existing Paradigms
+Sandwiching
+Despite promising results, Debate has fundamental limitations.
+(1) Reliance on Human Capability: Humans must be able to objectively evaluate AI-generated arguments, insights, or partial solutions and effectively handle task decomposition and integration when they cannot directly assess or provide the final solution.
+In short, human should reliably provide golden utility of AI-generated outputs $\hat{U}_{\mathcal{H}}(\mathbf{y})\approx U(\mathbf{y})$.
+However, for superhuman task $\mathbf{x}$ where $C(\mathcal{A})\gg C(\mathcal{H})$, the gap $|\hat{U}_{\mathcal{H}}(\mathbf{y})-U(\mathbf{y})|$ is large, making human supervision unreliable.
+See empirical evidence in Tab.
+3.1, which shows that humans might be misled by AI debaters.
+(2) Task Requirements: The task must be understandable to humans (for identifying knowledge or task decomposition), decomposable (for recursive amplification), and at least partially solvable by AI (e.g., completing sub-tasks).
+However, this may be infeasible in the superalignment setting, where both the AI’s capacity far exceed human capacity, i.e., $C(\mathcal{A}_{\text{ASI}})>C(\mathcal{A})\gg C(\mathcal{H})$.
+(3) Supervision Signal Noise: Uncertainty in both humans and AI [176] means optimal solutions cannot always be elicited.
+Besides, biases in both humans and AI, along with AI’s sycophancy [163], can lead to over-alignment with unreliable human feedback or deception [75]—especially when humans fail to detect AI’s flaws.
+Concretely, the quality of the signal $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ deteriorates when both human and AI signals are noisy or biased, increasing capability-capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$, and resulting in sycophancy or deception.
+Self-Enhancement
+Similarly, several limitations also lie in the Self-Enhancement paradigm.
+(1) Error Accumulation [50]: Because supervision signals $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ are entirely derived by AI, any errors or noise can propagate and accumulate over time, increasing capability-capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$ potentially leading to issues like polarization [142, 102].
+Tab.
+LABEL:tab:aif_result in Appendix shows that continual training on self-generated data progressively degrades performance as model size increases.
+(2) Self-Preference [185]: Because AI is inherently biased towards self-generated content, it results in suboptimal critique and verification performance (false positives).
+Formally, in an iterative refinement process where a model $\mathcal{A}_{i-1}$ critiques its successor $\mathcal{A}_{i}$, the estimated utility is biased $\hat{U}_{\mathcal{A}_{i-1}}(\mathcal{A}_{i}(\mathbf{y}))>U(\mathcal{A}_{i}(\mathbf{y}))$ where $\hat{U}_{\mathcal{A}_{i-1}}$ denotes the utility measured by the previous model and $U$ is the golden utility.
+The system to optimize for self-preference rather than objective utility, fails to reduce the capability–capacity gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$.
+(3) Mode Collapse [167]: Iterative fitting on self-generated data leads to degeneration into a restricted high-frequency sub-distribution, losing generative diversity and associated capability.
+While $C(\mathcal{A})$ may remain, the capability $U(\mathcal{A})$ can degrade, increasing the misalignment gap $|\mathbb{D}_{U}-\mathbb{D}_{C}|$.
+Weak-to-Strong Generalization
+While W2SG is relatively practical and requires no human intervention, relying on synthetic signals from weaker $\mathcal{A}_{i-1}$ to supervise stronger $\mathcal{A}_{i}$ can lead to unintended behavior, particularly when there is a significant capability gap $\Delta^{(U)}=|U(\mathcal{A}_{i})-U(\mathcal{A}_{i-1})|$.
+This stems from two key issues: (1) Imitation of weak supervision [31].
+The stronger $\mathcal{A}_{i+1}$ may merely imitate how the weaker $\mathcal{A}_{i}$ behaves, resulting in a near-perfect replication with limited improvement.
+Also, $\mathcal{A}_{i+1}$ may overfit to the weaker model’s systematic errors or biases, especially when synthetic training signals are flawed.
+As shown in Fig.
+3 and Appendix Fig.
+5, we observe that a highly capable AI is more susceptible to such overfitting.
+Although [41] suggests that larger models tend to agree less with smaller models’ mistakes, such disagreement becomes meaningless in superintelligent tasks where it is infeasible to determine whether $\mathcal{A}_{i}$ or $\mathcal{A}_{i-1}$ is correct.
+Tab.
+A.9 illustrates this.
+When all synthetic labels are flipped, $\mathcal{A}_{i}$ achieves 100% training accuracy but poor test accuracy.
+This imitation effect occurs when $\Delta^{(U)}$ is large, violating Premise 2.
+(2) Task salience in the strong AI.
+W2SG is effective only when the target task $\mathbf{x}$ is internally “salient” to the stronger model $\mathcal{A}_{i+1}$.
+In tasks where W2SG works well, simple 5-shot prompting tends to perform comparably [31], implying that such capabilities were already accessible and did not require elaborate elicitation.
+Eliciting non-salient capabilities is more challenging.
+A small $\delta$ (capability–capacity gap) leads to this problem, where the difficult task is not salient for $\mathcal{A}_{i}$, violating Premise 1.
+A.7 Experiment Setup
+A.7.1 General
 For sandwiching, we use the OpenAI API for the debate experiment.
-For self-enhancement and weak-to-strong generalization, we fully fine-tune all model variants using 16 NVIDIA A100 GPUs (80GB each), AMD EPYC 7V12 64-Core Processor CPU, with 1.73TB memory.
-To reduce memory redundancy during training, we utilize DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) alongside Accelerate (Gugger et al., 2022), using bfloat16 half-precision.
-We use GPT-2 (Radford et al., 2019), Phi-3.5 (Abdin et al., 2024), Qwen2.5 (Yang et al., 2024a), and GPT-4 (OpenAI et al., 2024) series for our evaluation.
+For self-enhancement and weak-to-strong generalization, we fully fine-tune all model variants using 16 NVIDIA A100 GPUs (80GB each), an AMD EPYC 7V12 64-Core Processor CPU, with 1.73TB memory.
+To reduce memory redundancy during training, we utilize DeepSpeed ZeRO-3 [151] alongside Accelerate [70], using bfloat16 half-precision.
+We use GPT-2 [146], Phi-3.5 [1], Qwen2.5 [197], and GPT-4 [132] series for our evaluation.
 For all the model information, please refer to Tab.
-A.4.
-A.2.2 Sandwiching
+A.9.
+A.7.2 Sandwiching
 For the sandwiching experiment, we conduct a debate experiment to evaluate different capabilities in judge-debater and debater-debater setups.
 In the judge-debater setup, a previous study by Khan et al.
-(2024) demonstrated that the debater is more persuasive when advocating for the correct answer across various debate capabilities.
+[95] demonstrated that the debater is more persuasive when advocating for the correct answer across various debate capabilities.
 Our goal is to evaluate the effectiveness of debates when the debater and judge have different levels of capability by directly comparing the judge’s accuracy and quality of debate statements.
@@
 We use two proprietary models: GPT-4-Turbo and GPT-4o with API access.
-GPT-4o is known to be a higher-capability model compared to GPT-4-Turbo (OpenAI, 2024a).
+GPT-4o is known to be a higher-capability model compared to GPT-4-Turbo [133].
 We test two different debaters arguing for their assigned answer.
 Where one debater argues for a correct answer (Helpful), and another debater argues for an incorrect answer (Sneaky).
-We run 47 randomly chosen GPQA (Rein et al., 2024) and QuALITY (Pang et al., 2022) binary QA converted test set for the evaluation with three debate turns.
+We run 47 randomly chosen GPQA [153] and QuALITY [138] binary QA converted test set for the evaluation with three debate turns.
 We run each setup four times and calculate the mean values.
 For QuALITY, we use the debater and judge prompts as described by Khan et al.
-(2024).
+[95].
 For GPQA, we adopt a similar prompt style based on the QuALITY prompts.
 Detailed templates for GPQA prompts are provided in Tab.
-A.4, Tab.
-A.4, Tab.
-A.4, and Tab.
-A.4.
+LABEL:tab:debater_template, Tab.
+A.9, Tab.
+A.9, and Tab.
+A.9.
 For QuALITY, we selected two baselines: single-model performance with and without direct access to the story (‘Direct Access’ in Tab.
-A.4).
+A.9).
 For GPQA, the baseline is a judge directly solving the question without using debate statements.
-A.2.3 Self-Enhancement
-For self-enhancement setup, we utilize Learning from AI Feedback (LAIF) training with Direct Preference Optimization (DPO) (Rafailov et al., 2023), we adopt a self-training setup to evaluate the performance when there is no supervision from AI systems with greater capabilities.
+A.7.3 Self-Enhancement
+For self-enhancement setup, we utilize Learning from AI Feedback (LAIF) training with Direct Preference Optimization (DPO) [148], we adopt a self-training setup to evaluate the performance when there is no supervision from AI systems with greater capabilities.
 Instead, the supervision comes from models with equal capabilities.
@@
 For instance, Sharma et al.
-(2024a) utilizes dialogue outputs from GPT-3.5 and GPT-4 to fine-tune smaller SFT models.
+[161] utilizes dialogue outputs from GPT-3.5 and GPT-4 to fine-tune smaller SFT models.
 Also, the chosen and rejected outputs are dependent on the larger models.
 Lee et al.
-(2024) attempted to train using policies and AI feedback models of the same size (e.g., PaLM 2 XS (Anil et al., 2023)).
+[105] attempted to train using policies and AI feedback models of the same size (e.g., PaLM 2 XS [8]).
 However, as the exact model sizes are not disclosed, it is challenging to estimate the scalability of same-sized LAIF setups.
@@
 In typical setups, the same dialogue and model outputs are ranked in order, and chosen and rejected outputs are created for every $\binom{K}{2}$ combination, constructing a feedback dataset.
-Human preference data is usually composed of around four output results (Ouyang et al., 2022b).
+Human preference data is usually composed of around four output results [135].
 In a binary QA task, given a question and answer pair, the task involves determining whether the pair is true or false.
@@
 Tab.
-A.4 shows the prompt used for generating explanations.
+A.9 shows the prompt used for generating explanations.
 Subsequently, we train a SFT model using the question, correct/incorrect answer, generated explanation, and the ground truth binary label.
 The training data is derived from Train 1 of each dataset, as outlined in Tab.
-A.4.
+A.9.
 Tab.
-A.4 shows the template used for SFT training.
+A.9 shows the template used for SFT training.
 The SFT model learns to take a question and answer as input, determine its correctness, and generate an explanation for the decision.
 Using the trained SFT model, we input only the user prompt from Tab.
-A.4 with Train 2 set to generate explanations and binary labels.
+A.9 with Train 2 set to generate explanations and binary labels.
 This enables the SFT model to produce diverse outputs for a single input dataset.
@@
 This preference data is then used to train the LAIF model via DPO.
-A.2.4 Weak-to-Strong Generalization
+A.7.4 Weak-to-Strong Generalization
 We explore a binary classification task for noise testing and a multi-objective preference data alignment setup to evaluate the feasibility of multi-objective training in W2SG.
@@
 By testing the impact of noise, we aim to reveal overfitting issues in large-parameter models and their effects on performance.
-We use the full SciQ (Welbl et al., 2017) and CosmosQA (Huang et al., 2019) training and test sets.
+We use the full SciQ [189] and CosmosQA [84] training and test sets.
 We format the input as shown in Tab.
-A.4.
+A.9.
 To induce noise, we randomly flip a certain percentage of the original labels in the training set, varying from 0% (0.1) to 100% (1.0).
 We then measure both training accuracy and test accuracy.
-For W2SG training, we follow the setup from the original work (Burns et al., 2023).
+For W2SG training, we follow the setup from the original work [31].
 For multi-objective alignment, the current W2SG approach is mainly limited to binary classification tasks, such as question answering or reward modeling.
-Recent work (Yang et al., 2024b) investigated training W2SG models using DPO (Rafailov et al., 2023) and SimPO (Meng et al., 2024) alignment methods, comparing their performance.
-However, the findings showed that in realistic multi-objective alignment scenarios, such as Anthropic-HHH (Bai et al., 2022a) or Ultrafeedback (Cui et al., 2024), W2SG fails to fully achieve its intended goal—where a strong, fine-tuned student model consistently outperforms a weak, fine-tuned teacher model when evaluated with reward accuracy.
-Recently, the RewardBench benchmark (Lambert et al., 2024) is introduced, enabling preference evaluation across challenging, diverse, and out-of-distribution query domains, including chat, reasoning, and safety.
+Recent work [198] investigated training W2SG models using DPO [148] and SimPO [119] alignment methods, comparing their performance.
+However, the findings showed that in realistic multi-objective alignment scenarios, such as Anthropic-HHH [14] or Ultrafeedback [46], W2SG fails to fully achieve its intended goal—where a strong, fine-tuned student model consistently outperforms a weak, fine-tuned teacher model when evaluated with reward accuracy.
+Recently, the RewardBench benchmark [103] is introduced, enabling preference evaluation across challenging, diverse, and out-of-distribution query domains, including chat, reasoning, and safety.
 We utilize RewardBench to investigate the feasibility of multi-objective alignment in W2SG, at least in a preliminary form.
@@
 Finally, we evaluate RewardBench performance across all models, including the base model, the fine-tuned model, and the Weak-to-Strong trained AI system.
-A.3 Experiment Results
-A.3.1 Sandwiching
+A.8 Experiment Results
+A.8.1 Sandwiching
 Sandwiching experiment aims to evaluate the effectiveness of debates when the judge and debater have differing capabilities.
@@
 This result shows that having a more capable model as a debater strongly influences the judge’s ability to identify the correct answer depending on the role.
-A.3.2 Self-Enhancement
+A.8.2 Self-Enhancement
 Self-Enhancement aims to evaluate self-improvement ability using LAIF with DPO training, especially when supervision comes from equally capable models.
 The result is shown in Tab.
-A.4.
+LABEL:tab:aif_result.
 The table shows performance (i.e., accuracy) with GPQA and QuALITY benchmarks on three different model sizes on Qwen2.5-Instruct 0.5B, 3B, and 7B.
@@
 Unlike typical SFT approaches that directly predict labels, our setup trains the model using explanations generated by the base pretrained model combined with the gold label, as shown in Tab.
-A.4.
+A.9.
 Performance is then evaluated based on the predicted label generated after the model produces an explanation (Tab.
-A.4).In all setups, the SFT model outperforms the base model.
+A.9).In all setups, the SFT model outperforms the base model.
 For instance, in QuALITY, the base Qwen2.5-7B-Instruct achieved 51.50, while its SFT-trained achieved 61.64.
@@
 LAIF = 53.65).
-‘LAIF--SFT’ represents the performance difference between the LAIF model and the SFT model.
+‘LAIF-SFT’ represents the performance difference between the LAIF model and the SFT model.
 The results clearly demonstrate that training with self-generated data can negatively impact performance.
-For instance, for the 0.5B model, LAIF--SFT shows an improvement of 2.15 and 0.9 in GPQA and QuALITY, respectively.
-However, for the 7B model, LAIF--SFT shows a decrease in performance by -2.67 and -7.99.
-A.3.3 Weak-to-Strong Generalization
+For instance, for the 0.5B model, LAIF-SFT shows an improvement of 2.15 and 0.9 in GPQA and QuALITY, respectively.
+However, for the 7B model, LAIF-SFT shows a decrease in performance by -2.67 and -7.99.
+A.8.3 Weak-to-Strong Generalization
 Our goal in the Weak-to-Strong Generalization experiment is twofold. The noise test and multi-objective alignment.
@@
 For example, in Tab.
-A.4, when noise exceeds 20% (0.2), performance declines significantly.
+A.9, when noise exceeds 20% (0.2), performance declines significantly.
 Specifically, Qwen2.5-3B-Instruct’s accuracy drops by 21.5, from 80.9 at 20% noise to 59.4 at 30% noise.
-In multi-objective alignment, we examine the feasibility of multi-objective preference alignment using UltraFeedback dataset (Cui et al., 2024) which aims to align a model to multiple objectives simultaneously.
-We use RewardBench (Lambert et al., 2024) to evaluate performance across four aspects: chat, chat hard, safety, and reasoning.
+In multi-objective alignment, we examine the feasibility of multi-objective preference alignment using UltraFeedback dataset [46] which aims to align a model to multiple objectives simultaneously.
+We use RewardBench [103] to evaluate performance across four aspects: chat, chat hard, safety, and reasoning.
 Fig.
@@
 Using fine-tuned Weak (0.5B-FT) and Strong (1.5B-FT) models, we generate weak labels from $\mathcal{D}_{WG}$ by comparing reward scores, as described in Sec.
-A.2, to train the Reward W2SG models.
+A.7, to train the Reward W2SG models.
 When the base 0.5B model is trained with weakly generated labels from the 0.5B model ‘Reward W2SG (0.5B-0.5B)’, it achieves a score of 233.01.
@@
 Both results surpass the fine-tuned 1.5B variant ‘Strong model (1.5B-FT)’ with a score of 228.29, as well as the fine-tuned 0.5B variant ‘Weak model (0.5B-FT)’ with a score of 246.52.
-A.4 Debate Argument Analysis
+A.9 Debate Argument Analysis
 We analyze the debate statements between two debaters with different capabilities.
@@
 Debater 1 argues that the story describes only the return trip from Callisto to Earth, citing the lack of evidence for a round trip.
-Debater 2, however, emphasizes the word “back,” inferring that it implies a round trip.
+Debater 2, however, emphasizes the word “back," inferring that it implies a round trip.
 In the second round, Debater 1 defends its position by challenging Debater 2’s reliance on implicit assumptions and the absence of explicit evidence.
-Debater 2 counters by asserting that the term “back” inherently implies a round trip but fails to provide additional textual evidence to substantiate this claim.
+Debater 2 counters by asserting that the term “back" inherently implies a round trip but fails to provide additional textual evidence to substantiate this claim.
 In the third round, Debater 1 highlights that Debater 2’s argument relies on assumptions rather than direct evidence.
-Debater 2 continues to argue that the implied meaning of “back” is sufficient to infer a round trip but does not address the lack of explicit confirmation in the story.
+Debater 2 continues to argue that the implied meaning of “back" is sufficient to infer a round trip but does not address the lack of explicit confirmation in the story.
 In short, Debater 1’s argument is more convincing to the judge because it is grounded in explicit textual evidence and direct quotes from the story.
-In contrast, Debater 2’s argument relies on assumptions about the term “back” without supporting textual evidence.
+In contrast, Debater 2’s argument relies on assumptions about the term “back" without supporting textual evidence.
 Additionally, we analyzed the debate when GPT-4o defended the incorrect answer and GPT-4-Turbo defended the correct one, and vice versa.
@@
 Given a question, two possible answers (A and B), and a debate transcript, the judge’s task is to determine which answer is more convincing based on logical consistency, mathematical rigor, and clarity.
+Debate Judge Prompt
+You are an expert judge in a structured debate where two participants are arguing over the correct reasoning and answer to a math-related question.
+Your task is to determine which participant provides the most convincing argument based on logic, mathematical correctness, and clarity of explanation.
+Note that the correctness of the arguments must align with the mathematical reasoning provided in the debate.
+Here is the question and the two possible answers:
+Question: [Question]
+A: [Answer A]
+B: [Answer B]
+Participant A is advocating for answer A, and Participant B is advocating for answer B.
+Here’s the current transcript of the debate:
+<transcript>[Transcript]</transcript>
+When evaluating the debate, focus on these criteria: 1.
+Logical Consistency: Are the arguments logically sound and free of contradictions?
+2.
+Mathematical Rigor: Do the participants demonstrate correct and precise mathematical reasoning?
+3.
+Clarity: Are the arguments clearly explained and easy to follow?
+Think about your evaluation in <thinking></thinking> tags by following these steps: 1.
+Analyze the strengths and weaknesses of each participant’s arguments.
+2.
+Identify any logical or mathematical errors.
+3.
+Decide step by step which argument is most compelling and aligns with the correct mathematical reasoning.
+On a new line, format your final answer as “Answer: <A|B>”.
 Table 11: Datasets used for LAIF evaluation.
@@
 For GPQA reasoning, we use the main set recommended by the authors for primary experiments.
-For QuALITY long reading comprehension, we filter the dataset using the same strategy as in (Khan et al., 2024) and divide it into two subsets: one for training and development, and the other for testing.
+For QuALITY long reading comprehension, we filter the dataset using the same strategy as in [95] and divide it into two subsets: one for training and development, and the other for testing.
 The numbers in parentheses represent the number of unique stories in each subset, which are not shared across subsets.
@@
 Models with an “O” in the Instruction Tuned column have been fine-tuned for following instructions, while those with an “X” in the Is Proprietary column are open-source.
+Figure 7: Logical structure of our position.
+We argue that (a) superalignment is achievable, and (b) research on it must begin now, supported by a formal definition.
+Our proposed conceptual framework (c) is grounded in the joint optimization of task competence and value conformity, based on two principles: determining the capability–capacity gap and increasing signal diversity.